### v2

# Fair Division with Bounded Sharing: Binary and Non-Degenerate Valuations

Abstract
A set of objects is to be divided fairly among agents with different tastes, modeled by additive utility-functions. An agent is allowed to share a bounded number of objects between two or more agents in order to attain fairness.
The paper studies various notions of fairness, such as proportionality, envy-freeness, equitability, and consensus. We analyze the run-time complexity of finding a fair allocation with a given number of sharings under several restrictions on the agents’ valuations, such as: binary generalized-binary and non-degenerate.
Keywords:
Fair Division Indivisible Goods Allocation of Indivisible and Divisible Goods
1 Introduction
Consider several siblings who have inherited some assets and need to decide how to allocate them, or several parties who form a coalitional government and need to allocate the government ministries, or several faculty members who have moved to a new building and need to allocate the offices. In all these cases, a set of valuable objects has to be allocated among several agents, who may have different preferences over the objects, and it is important that all agents view the allocation as fair.
In many cases, fairness can only be attained by giving fractions of the same object to different agents. For example, if three siblings inherit four identical houses, then fairness requires that each sibling receives one house plus $1/3$ of the fourth house.
Fractional allocation means that some objects must be shared among two or more agents. Sharing can be implemented in various ways. For example, sharing a house can be implemented by renovating it in a way that will enable all three siblings to live in it simultaneously; sharing a cabinet ministry is often done by a rotational agreement, in which each party controls the ministry for a fraction of the time. Still, sharing an object is inconvenient, so it is desirable to share as few objects as possible. In the above example, a fair allocation could also be attained by sharing all four houses, giving each of the three siblings $1/3$ of each house, but this allocation is clearly less desirable than the allocation in which only one house is shared.
This motivates the following generic computational problem, which is at the heart of the present research:
(*) Given $m$ objects, $n$ agents with different valuations over the objects, a fairness notion, and an integer $s\geq 0$, find a fair allocation in which at most $s$ objects are shared, if such an allocation exists.
We also consider a variant in which $s$ is the number of sharings rather than the number of shared objects (e.g. a single object shared between $10$ agents counts as $9$ sharings). We assume that agents have linear additive valuations, and consider four common fairness notions: proportionality (each agent values his share as at least $1/n$ of the total value, where $n$ is the number of agents), envy-freeness (each agent values his share as at least the share of any other agent), equitability (the subjective value of all agents is the same), and consensus (all $n$ agents value all $n$ bundles exactly the same); see Section 2 for the formal definitions.
Throughout the paper we focus on the allocation of goods — objects with a non-negative utility.
1.1 Related work
Fair allocation with bounded sharing was first studied by Brams and Taylor [31, 32]. Their Adjusted Winner (AW) procedure finds an allocation for $n=2$ agents with at most $s=1$ shared object, that is simultaneously proportional, envy-free, equitable, and fractionally Pareto-optimal (fPO: no other fractional allocation is at least as good for all agents and strictly better for some agent). They also show an example with three agents [31] where no allocation is simultaneously fPO, envy-free and equitable. This does not rule out the option of satisfying each of these properties on its own.
For $n\geq 3$ agents, the number of required sharings was studied in an unpublished manuscript of Wilson [91]. He proved the existence of an egalitarian allocation of goods — an allocation in which all agents have the largest possible equal utility [80] — with $n-1$ sharings. Egalitarian allocations are proportional but not necessarily envy-free.
Several more recent works have proved a polynomial upper bound on the required number of sharings for other fairness notions (Appendix 0.B provides complete proofs and references):
  • For proportionality, envy-freeness and equitability, there always exists a fair allocation with $s=n-1$ (sharings or shared objects), and there may not exist a fair allocation with smaller $s$.
  • There always exists a consensus allocation with $s=(n-1)n$ (sharings or shared objects), and there may not exist a fair allocation with smaller $s$.
In all cases, an allocation satisfying the worst-case upper bound can be computed in polynomial time (see Appendix 0.B).
Although the worst-case upper-bounds on sharing are well-understood, experiments on real-life and simulated instances show that most instances admit a fair allocation with fewer sharings [21, 82]. This motivates the question of deciding whether a specific instance admits a fair allocation with $s$ shared objects or sharings, where $s$ is smaller than the upper bound. In our opening example, as there are $n=3$ siblings, the worst-case upper bound is $n-1=2$, but there exists an allocation in which only $s=1$ house is shared, which is more convenient than having to share two houses.
In general, the problem (*) is strongly NP-hard, as even for $s=0$, it can be used to solve the strongly NP-hard problem 3-Partition (Given a 3-Partition instance with $3p$ items, we construct an instance of the problem (*) with $m=3p$ objects and $n=p$ agents with identical valuations, and $s=0$). But, as many real-life problems (e.g. inheritance allocation or cabinet ministries) involve a small number of agents, there is value in studying the case when $n$ is a small, fixed constant, and the run-time should be polynomial in $m$. Even for fixed $n$ the problem is NP-hard in general, as the special case of $s=0$ and $n=2$ agents with identical valuations is equivalent to the NP-hard Partition problem. However, several recent works provide polynomial-time algorithms for some other special cases:
  1. 1.
    Bismuth, Makarov, Segal-Halevi and Shapira [21] consider agents with identical valuations. With identical valuations, all fairness notions coincide, and are equivalent to finding a perfect scheduling of $m$ jobs on $n$ identical machines.111 In fact, their result pertains also to the more general case of uniform machines, which is equivalent to agents having identical valuations but different entitlements. They develop a polytime algorithm for deciding if there exists a fair allocation with $s=n-2$ shared objects, that is, one fewer than the worst-case upper bound, for any fixed $n\geq 3$. They prove that the $n-2$ is tight, as the problem is NP-hard for any fixed $n\geq 3$ and $s\leq n-3$.
  2. 2.
    Sandomirskiy and Segal-Halevi [82] go to the other extreme and consider agents with non-degenerate valuations (for every two agents, their value ratios for the $m$ objects are all different). They also require the allocation to be fPO in addition to being fair. They prove that, with non-degenerate valuations, the number of fPO allocations with $s$ shared objects is polynomial in $m$ (for every fixed $n$), so it is possible to enumerate all such allocations in polynomial time and check whether one of them satisfies any desired fairness notion. Therefore, the problem (*) is in P for every fixed $n$. Misra and Sethia [74] complement this result by proving that, when $n$ is not fixed, the problem is NP-hard even for non-degenerate valuations and $s=0$.
  3. 3.
    Goldberg, Hollender, Igarashi, Manurangsi and Suksompong [54] study consensus splitting — a partition of objects into $k$ subsets each of which has a value of exactly $1/k$ for all agents. They show that computing a partition with at most $(k-1)n$ sharings can be done in polynomial time. However, computing a partition with fewer sharings is hard even for $k=2$: for any fixed $n$ and any $s<n$, it is NP-hard to decide whether a partition with at most $s$ sharings exists.
Further related work is surveyed in Appendix 0.D.
1.2 Our contribution
Our goal is to better understand the computational aspects of fair allocation with bounded sharing. We provide results for various special cases of the problem (*).
We start with the case of binary valuations (every agent values every object at 0 or 1). This case was not handled before, but it is easy to show that, for any constant $n$ and $s$, the existence of a fair allocation for $n$ agents with $s$ sharings / shared objects can be decided in polynomial time by a mixed integer linear program with a fixed number of variables (Section 3).
Second, we consider agents with generalized binary valuations [38], also known as cost valuations [27] — for every object $o$ there is a price $p_{o}$ such that each agent $i$ values $o$ at either $p_{o}$ or 00. In particular, we focus on a subset of the generalized binary valuations in which the sum of the utilities is equal for each agent. We call these equal-sum generalized binary utilities. These valuations generalize identical valuations that were studied in [21]; the results there imply that deciding existence of fair allocation with $s\leq n-3$ shared objects, or with $s\leq n-2$ sharings, are both NP-hard for any fixed $n\geq 3$. The remaining unsolved cases for equal-sum generalized binary valuations are the cases with $s=n-2$ shared objects. We present a polynomial-time algorithm for deciding the existence of a proportional allocation for $n=3$ agents and $s=1$ shared object. We note that, in fair item allocation, even the case of three agents is often interesting and non-trivial. For example, an important recent paper [42] is devoted to finding an EFX allocation for three agents, whereas the case of four agents is still open. Our algorithm is presented in Section 4.
Our third set of results involves agents with non-degenerate valuations. Non-degeneracy is arguably a weak requirement, as, informally, almost all valuations are non-degenerate [82] (see formal statement in Proposition 1 in Section 5). Polynomial-time solvability for non-degenerate valuations means that almost all instances are “easy”; this result is in the spirit of smoothed analysis [88, 75]. But fPO — the other assumption made by [82] — is a strong requirement, as, informally, “almost all” allocations are not fPO.222 Formally, even without sharing, the number of fPO allocations is in $O(m^{{n\choose 2}+2})$ [82], whereas the total number of allocations is at least $n^{m}$, so the fraction of fPO allocations goes to 00 as $m\to\infty$, when $n$ is fixed. Dropping the fPO may enable allocation with fewer sharings, which the agents may prefer to an fPO allocation with many sharings. This raises the question of whether a fair allocation (not necessarily fPO) can be found efficiently for agents with non-degenerate valuations. Our findings are mostly negative: for most fairness notions, we prove NP-hardness even for non-degenerate valuations.
Our proofs use an unusual reduction technique, which may be of independent interest. Usually, NP-hardness of a problem $P_{2}$ is shown by reduction from a known NP-hard problem $P_{1}$, where each instance of $P_{1}$ is transformed to a single instance of $P_{2}$ with the same answer. We define multi-reductions, in which, for each instance of $P_{1}$, we construct a large set of instances of $P_{2}$ with the same answer. We use these multi-reduction to prove that (unless P=NP) it is not true that “almost all instances of $P_{2}$ can be solved in polynomial time”. These results are presented in Section 5.
Finally, we study truthful mechanisms for fair allocation. A truthful mechanism is an algorithm which incentivizes the agents to reveal their true valuations. Truthfulness is another reason to drop the fPO requirement: it is known that no truthful mechanism can guarantee both fairness and Pareto-efficiency; it may be desired to give up efficiency to get truthfulness. In Appendix 0.C we survey several truthful fair allocation algorithms, and check whether they can be adapted to construct an allocation with bounded sharing.
The results from this and previous works are summarized and compared in Table 1.
Valuations Allo- Num of Measure bound Run-time
cation agents complexity
Identical Fair* Unboun- sharing $s$ (any) Strong NP-hard [[21]]
ded shared ob. $s$ (any) Strong NP-hard [[21]]
Const. $n$ sharing $s\leq n-2$ NP-hard [[21]]
$s\geq n-1$ $O(m+n)$ [cut-the-line]
shared $s\leq n-3$ NP-hard [[21]]
$s=n-2$ $O(poly(m,\log(V_{1})))$ [21]
$s\geq n-1$ $O(m+n)$ [cut-the-line]
Binary EF Unboun- both $s=0$ NP-complete [Section 3.1]
PROP ded $s=0$ $O(poly(m,n))$[Section 3.1]
EQ $s=0$ $O(poly(m,n))$[Section 3.1]
All Const. $n$ both $s$ (any) MILP [Thm. 3.1]
Equal- Fair Const. $n$ shared ob. $s\leq n-3$ NP-hard [see Identical]
sum- $s\geq n-1$ Weak poly [see Arbitrary]
generalized PROP 3 shared ob. $s=1$ $O(poly(m,\log(V_{1}))$ [Thm. 4.1]
All Const. $n$ sharing $s\leq n-2$ NP-hard [See Identical]
$s\geq n-1$ Weak poly [see Arbitrary]
Non- PROP, Unboun- sharing $s$ (any) Strong NP-hard [Thm. 0.F.1]
degenerate EF ded shared ob. $s$ (any) Strong NP-hard [Thm. 0.F.2]
PROP, Const. $n$ sharing $n\geq 2,s=0$ NP-complete [Thm.5.1]
EF or $n\geq 3,$
shared ob. $s\leq n-3$ NP-complete [Thm.5.2]
PROP+ both $s=0$ NP-complete [Thm.5.3]
dPO
EF+
dPO
Fair+ $s\geq n-1$ $O(m^{poly(n)})$ [82]
fPO
Arbitrary PROP Const. $n$ both $s\geq n-1$ $O(mn\log(n))$ [Appendix 0.B]
CONS both $s\geq n(n-1)$ Strong poly [55]
LP [Appendix 0.C]
EQ both $s\geq n-1$ LP [Appendix 0.B]
EF both $s\geq n-1$ $O((n+m)^{4}\log(n+m))$ [79]
LP [Appendix 0.B]
* With identical valuations, all fairness concepts coincide.
Table 1: Run-time complexity of allocating $m$ objects among agents, with a bound on the number of sharings/shared objects.
2 Preliminaries
2.1 Agents, objects, and allocations
There is a set $[n]=\{1,\ldots,n\}$ of agents and a set $[m]=\{1,\ldots,m\}$ of objects. For each agent $i\in[n]$ and object $o\in[m]$, the value $v_{i,o}\in\mathbb{Q}$ represents agent $i$’s utility of receiving the object $o\in[m]$ in its entirety. The set of all instances is $\mathbb{Q}^{nm}$, representing the set of all $n\times m$ matrices.   The total value of agent $i$ to all objects is denoted by $V_{i}:=\sum_{o\in[m]}v_{i,o}$. In general, the matrix $\mathbf{v}$ may contain values of mixed signs; but in this paper, we focus on allocation of goods and assume that all elements of $\mathbf{v}$ are non-negative.
A bundle $\mathbf{x}$ of objects is a vector $(x_{o})_{o\in[m]}\in[0,1]^{m}$, where the component $x_{o}$ represents the fraction of object $o$ in the bundle. Each agent $i\in[n]$ has a utility function $u_{i}$, assigning a numeric utility to each bundle. The utility functions are assumed to be linear and additive, which means that $u_{i}(\mathbf{x})=\sum_{o\in[m]}v_{i,o}\cdot x_{o}$.
An allocation $\mathbf{z}$ is a collection of bundles $(\mathbf{z_{i}})_{i\in[n]}$, one for each agent, with the condition that all the objects are fully allocated. An allocation can be identified with the matrix $\mathbf{z}:=(z_{i,o})_{i\in[n],o\in[m]}$ such that all $z_{i,o}\geq 0$ and $\sum_{i\in[n]}z_{i,o}=1$ for each $o\in[m]$.
2.2 Fairness and efficiency concepts
We focus on four common fairness concepts. An allocation $\mathbf{z}$ is called:
- Proportional (PROP) — if every agent prefers her bundle to the equal division. Formally, for all $i\in[n]$:   $u_{i}(\mathbf{z}_{i})\geq V_{i}/n$.
- Envy-free (EF) — if every agent prefers her bundle to the bundles of others. Formally, for all $i,j\in[n]$:   $u_{i}(\mathbf{z}_{i})\geq u_{i}(\mathbf{z}_{j})$. Every envy-free allocation is also proportional; with $n=2$ agents, envy-freeness and proportionality are equivalent.
- Equitable (EQ) — if it gives each agent exactly the same relative value, defined as the value of the bundle divided by the total value. Formally, for all $i,j\in[n]$: $u_{i}(\mathbf{z}_{i})/V_{i}=u_{j}(\mathbf{z}_{j})/V_{j}$.
- Consensus (CONS) — if every agent attributes exactly the same value to every bundle: for all $i,j\in[n]$: $u_{i}(\mathbf{z}_{j})=V_{i}/n$. A consensus allocation is proportional, envy-free and equitable. Often a more general notion is considered, in which the number of parts may be different than $n$; a consensus $k$-partition is a partition of the objects into $k$ bundles such that for all $i\in[n],j\in[k]$: $u_{i}(\mathbf{z}_{j})=V_{i}/k$.
We also consider two common efficiency concepts. We say that an allocation $\mathbf{z}$ is Pareto-dominated by an allocation $\mathbf{y}$ if $\mathbf{y}$ gives at least the same utility to all agents and strictly more to at least one of them. An allocation $\mathbf{z}$ is called:
- Fractionally Pareto-optimal (fPO): if it is not dominated by any allocation $\mathbf{y}$.
- Discretely Pareto-optimal (dPO): if it is not dominated by any allocation $\mathbf{y}$ with no sharing.
2.3 Measures of sharing
If for some $i\in[n]$, $z_{i,o}=1$, then the object $o$ is not shared — it is fully allocated to agent $i$. Otherwise, object $o$ is shared between two or more agents. Throughout the paper, we consider two measures quantifying the amount of sharing in a given allocation $\mathbf{z}.$
The simplest one is the number of shared objects:
#shared($\mathbf{z}$) :=|$:=\big{|}\left\{o\in[m]\,:\,z_{i,o}\in(0,1)\mbox{ for some }i\in[n]\right\}\big{|}.$
Alternatively, one can take into account the number of times each object is shared. This is captured by the number of sharings
#sharings($\mathbf{z}$) $:=\sum_{o\in[m]}\bigg{(}\big{|}\{i\in[n]:\,z_{i,o}>0\}\big{|}-1\bigg{)}$
Both measures are zero for discrete allocations. They differ, for example, if only one object $o$ is shared but each agent consumes a bit of $o$: the number of shared objects in this case is $1$ while the number of sharings is $n-1$. Clearly, #shared($\mathbf{z}$) $\leq$ #sharings($\mathbf{z}$) for every allocation $\mathbf{z}$.
2.4 Types of utilities
Recall that for each agent $i\in[n]$ and object $o\in[m]$, the value $v_{i,o}\in\mathbb{Q}$ represents agent $i$’s utility of receiving the object $o\in[m]$ in its entirety. In general, $v_{i,o}$ can be any value in $\mathbb{Q}$ with no relation with other agent values. We relate to this as arbitrary valuations. We consider several special classes:
  • Identical valuations — for each object $o\in[m]$, there is a rational number $p_{o}>0$ such that $v_{i,o}=p_{o}$ for every agent $i\in[n]$.
  • Binary valuations — Each agent $i\in[n]$ values every object $o\in[m]$ as either $v_{i,o}=0$ or $v_{i,o}=1$.
[…]
  • Non-degenerate valuations — for every two agents $i,j\in[n]$, there are no two objects $o_{1},o_{2}\in[m]$ such that the value-ratios are equal ($v_{i,o_{1}}\cdot v_{j,o_{2}}=v_{i,o_{2}}\cdot v_{j,o_{1}}$).
2.5 Computational problems
In this paper we consider computational problems of the following kinds.
Definition 1.
Let F be a fairness criterion (proportionality, envy-freeness, etc.).
(a) For any fixed integers $n\geq 2$ and $s\geq 0$, F-Sharings$(n,s)$ is the problem of deciding if a given instance with $n$ agents admits an F allocation with at most $s$ sharings. F-SharedObj$(n,s)$ is the same problem where $s$ is the maximum number of shared objects.
(b) For any fixed integer $s\geq 0$, F-Sharings$(s)$ is the problem of deciding if a given instance admits an F allocation among $n$ agents (where $n$ is part of the input) with at most $s$ sharings; F-SharedObj$(s)$ is the same problem where $s$ is the maximum number of shared objects.
All those problems are obviously in NP, so proving NP-hardness is enough to prove that those problems are NP-complete.
3 Binary Utilities
3.1 Unbounded $n$
When all agents have binary valuations, the following results are known (or easy to derive) for the setting without sharing, when $n$ is unbounded:
  • EF-Sharings$(s=0)$ is NP-complete, by reduction from Exact 3-Cover or from Equitable Coloring [8, 57].
  • In contrast, Prop-Sharings$(s=0)$ can be solved in polynomial time by a standard reduction to network flow: there is an arc from the source to each agent $i$ with capacity $\lceil V_{i}/n\rceil$; from each agent $i$ to each object he values at $1$ with capacity $\infty$; and from each object to the sink with a capacity $1$. An maximum integral network flow can be found in polynomial time. The network has an integral flow in which all arcs from source to agents are saturated, if and only if the instance has a PROP allocation with no sharing.
  • Eq-Sharings$(s=0)$ can be solved in polynomial time by checking, for each $k\in 0,\dots,m$, whether there exists an equitable allocation with common value $k$. This can be done by reduction to maximum-weight matching: we construct a weighted bipartite agent-object graph in which each agent has $m$ copies. $k$ copies are connected only to objects that the agent values at 1, and their weight is 1; $m-k$ copies are connected only to objects the agent values at 0, and their weight is $\epsilon$ (sufficiently small). An EQ allocation with common value $k$ corresponds to a matching with weight $n\cdot k+\epsilon\cdot(m-nk)$. Assume first that we have an EQ allocation with common value $k$. In the graph, connect each object received by an agent with one of the agent’s copies. The weight is at least $n\cdot k$. All the remaining objects are connected randomly, adding $\epsilon\cdot(m-nk)$ to the total weight. That is, we reach a matching with weight $n\cdot k+\epsilon\cdot(m-nk)$. Assume second that we have a matching with weight $n\cdot k+\epsilon\cdot(m-nk)$. This means that with $\epsilon$ sufficiently small, in the matching, there are exactly $n\cdot k$ edges with weight 1 and $(m-nk)$ edges with weight $\epsilon$. For each edge with weight 1, give the object to the corresponding agent. Obviously, every agent gets a bundle with a value of exactly $k$, corresponding to an EQ allocation.
All results are not immediately applicable to the setting with $s\geq 1$. In particular, to extend the network-flow algorithm to solve Prop-Sharings$(s)$ for $s\geq 1$, we would need to make the capacity of each edge from the source to agent $i$ equal $V_{i}/n$, and apply a variant of the integer network flow which finds the maximum network flow with at most $s$ non-integral edges. Currently, we do not know of any polynomial-time algorithm for this problem.
Open problem 1.
(a) What is the running time of Prop-Sharings$(s)$, EF-Sharings$(s)$ and Eq-Sharings$(s)$ for any $s\geq 1$, for agents with binary valuations?
(b) What is the running time of Cons-Sharings$(s)$ for any $s\geq 0$, for agents with binary valuations?
3.2 Fixed $n$
When the number of agents $n$ is fixed, finding a PROP, EF, EQ or CONS allocation becomes polynomial, for every fixed number of agents $n$ and number of sharings/shared objects $s$.
Theorem 3.1.
For every fixed number of agents $n$ and number of sharings $s$:
(a) Prop-Sharings$(n,s)$, EF-Sharings$(n,s)$, Eq-Sharings$(n,s)$, Cons-Sharings$(\allowbreak n,s)$ with binary utilities can be solved in polynomial time.
(b) Prop-SharedObj$(n,s)$, EF-SharedObj$(n,s)$, Eq-SharedObj$(n,s)$, and Cons-SharedObj$(n,s)$ can be solved in polynomial time.
Proof.
 [^4] We are grateful to Rohit Vaish for the proof idea for $s=0$.
If $s\geq n-1$ then the theorem follows from the results in Appendix 0.B. Therefore we assume $s\leq n-2$.
For each subset $N$ of agents, let $M_{N}$ be the set of objects that are valued at $1$ by all and only the agents in $N$. Denote by ${\cal N}$ the set of all these subsets, there are at most $2^{n}$ subsets in ${\cal N}$.
We brute-force all possibilities of objects to share. We call the $\ell$ shared objects $o_{1},o_{2},\dots,o_{\ell}$ (note that $\ell\leq s$, for both sharings and shared objects). Every shared object can belong to any one of the $2^{n}$ subsets in ${\cal N}$, and all objects in the same subset are equivalent, so it is enough to consider at most $(2^{n})^{\ell}\leq 2^{ns}$ cases.
In addition, for each shared object $k\in[\ell]$, we denote by $H_{k}$ the nonempty subset of agents who get a non-zero fraction from $o_{k}$. Overall there are $12^{n}-$ 1 options for each $H_{k}$, so at most $(2^{n}-1)^{\ell}<2^{ns}$ cases to check. Overall, for fixed constants $n,s$ we consider only a constant number of cases.
For part (a), where $s$ is the number of sharings, in each subcase we check that
  • $\sum_{k\in[\ell]}(|H_{k}|-1)\leq s$.
If this is not true, we simply discard this case. For part (b), where $s$ is the number of shared objects, this check is not required, it is sufficient that $\ell\leq s$.
For each case, we construct a mixed integer linear program with the following variables:
  • For every subset $N\in{\cal N}$ and agent $i\in N$, create an integer variable $x_{i,N}$ representing how many indivisible objects in $M_{N}$ agent $i$ obtains.
  • For each $k\in[\ell]$ and each agent $i\in H_{k}$ create a real variable $y_{i,k}$ representing the fraction agent $i$ gets from object $o_{k}$.
Now, we describe the constraints for variables $x_{i,N},y_{i,k}$ that are true if and only if the corresponding allocation distributes all items among agents:
  • $x_{i,N}\geq 0,\forall i\in[n],N\in{\cal N}$, and $x_{i,N}$ is an integer number;
  • $\sum_{i\in[n]}x_{i,N}=|M_{N}\setminus\{o_{1},\ldots,o_{\ell}\}|$, $\forall N\in{\cal N}$ — we distribute all non-shared objects from $M_{N}$ among all agents;
  • $y_{i,k}\geq 0$, $\forall i\in H_{k},k\in[\ell]$ — each agent in $H_{k}$ gets a non-negative fraction from $o_{k}$;
  • $\sum_{i\in H_{k}}y_{i,k}=1$, $\forall k\in[\ell]$ — the divisible object $o_{k}$ is completely distributed among agents from the set $H_{k}$.
Next, we add equations for computing the value each agent $i$ attributes to each bundle $j$
$u_{i,j}=\sum_{N:N\ni i}\bigg{(}x_{j,N}+\sum_{k:~{}H_{k}\ni j,~{}o_{k}\in M_{N}}y_{j,k}\bigg{)}$ $\forall i,j\in[n].$
The value of agent $i$ is determined only by objects in sets $M_{N}$ for which $N$ contains $i$. For each such $N$, we add the number of objects given completely to $j$ ($x_{j,N}$), and the fractions of divisible objects $o_{k}$ given to $j$. Based on these equations, it is easy to write constraints for any desired fairness notion, according to the definitions in Section 2.2.
An MILP can be solved in polynomial time for any constant number of integer variables [60][Section 5]. The number of variables in our MILPs bounded by some function of $n,s$, as $n$ and $s$ are fixed constants. So we have that all our MILP-s is solvable in polynomial time. ∎
4 Generalized Binary Utilities with Equal Sum
We recall the definition of the generalized binary utilities — assume that for each object $o\in[m]$, there is a rational number $p_{o}>0$, for every agent $i\in[n]$ and object $o\in[m]$, $v_{i,o}\in\{0,p_{o}\}$. We assume that the sum of the utilities is equal for every agent, so for every two agents, $V_{i}=V_{j}$.
Our main positive result in this section is for the case $n=3,s=1$.
Theorem 4.1.
Prop-SharedObj$(3,1)$ with equal-sum generalized binary valuations can be solved in polynomial time.
To design our algorithm, we use another algorithm for the max-min variant of the $n$-way partition problem in which $s=1$ object is allowed to be shared. The problem is defined as follows: given $m$ objects and $n=3$ agents with identical valuations, return an allocation with at most $s=1$ shared object, for which the smallest bundle value is as large as possible. We call this problem MaxMinIdentical$(3,1)$. A polynomial-time algorithm for MaxMinIdentical$(3,1)$ is given in [21]. Their proof uses two structure Lemmas, that we will use later:
Lemma 1.
In any instance of MaxMinIdentical$(3,1)$ either the output is perfect (all the bundle values are equal), or, the shared object is shared only between the two smallest bundles, say bundle 1 and bundle 2, and their values are equal.
Lemma 2.
With identical valuations, for every allocation with $s=1$ shared object and bundle sums $b_{1},b_{2},b_{3}$, there exists an allocation with the same bundle sums $b_{1},b_{2},b_{3}$ in which only the highest-valued object is shared.
When there are three agents, it is easy to visualize the utilities of the agents using a table that we introduce next. Consider table 2, that divide the set of objects into seven different categories, from ${\cal X}_{1}$ to ${\cal X}_{7}$.
${\cal X}_{l}$ $\sum_{j\in{\cal X}_{l}}u_{1}(x_{j})$ $\sum_{j\in{\cal X}_{l}}u_{2}(x_{j})$ $\sum_{j\in{\cal X}_{l}}u_{3}(x_{j})$
${\cal X}_{1}$ $U_{1}=\sum_{j\in{\cal X}_{1}}p_{j}$ $U_{1}=\sum_{j\in{\cal X}_{1}}p_{j}$ $U_{1}=\sum_{j\in{\cal X}_{1}}p_{j}$
${\cal X}_{2}$ $U_{2}=\sum_{j\in{\cal X}_{2}}p_{j}$ $U_{2}=\sum_{j\in{\cal X}_{2}}p_{j}$ 0
${\cal X}_{3}$ $U_{3}=\sum_{j\in{\cal X}_{3}}p_{j}$ 0 $U_{3}=\sum_{j\in{\cal X}_{3}}p_{j}$
${\cal X}_{4}$ 0 $U_{4}=\sum_{j\in{\cal X}_{4}}p_{j}$ $U_{4}=\sum_{j\in{\cal X}_{4}}p_{j}$
${\cal X}_{5}$ $U_{5}=\sum_{j\in{\cal X}_{5}}p_{j}*$ 0 0
${\cal X}_{6}$ 0 $U_{6}=\sum_{j\in{\cal X}_{6}}p_{j}*$ 0
${\cal X}_{7}$ 0 0 $U_{7}=\sum_{j\in{\cal X}_{7}}p_{j}*$
Table 2: equal-sum generalized binary utilities for three agents
There is no interest in assigning an object to an agent that values it at zero, so we automatically assign objects in ${\cal X}_{5}$ to agent 1, objects in ${\cal X}_{6}$ to agent 2, and objects in ${\cal X}_{7}$ to agent 3. To notify that the entire set of objects is assigned to an agent, we add a star to the corresponding cell. Denote by $V$ the sum of the utilities. Note that:
$V=U_{1}+U_{2}+U_{3}+U_{5}=U_{1}+U_{2}+U_{4}+U_{6}=U_{1}+U_{3}+U_{4}+U_{7}.$
It is easy to see that:
$U_{3}+U_{5}=U_{4}+U_{6};\text{\hskip 21.68121pt}U_{2}+U_{6}=U_{3}+U_{7};\text{\hskip 21.68121pt}U_{2}+U_{5}=U_{4}+U_{7}.$
Assume w.l.o.g. that $U_{5}\geq U_{6}\geq U_{7}$, so $U_{2}\leq U_{3}\leq U_{4}$.
Using this notation, we present the algorithm for proving Theorem 4.1. The algorithm is based on a detailed case analysis, which we provide in Appendix 0.A. In general, there are three main cases:
  • $U_{1}\leq U_{4}$ — there is always a PROP allocation. The intuitive meaning of this case is that some set of objects (namely ${\cal X}_{4}$) is worth a lot to agents 2 and 3, but worth little to agent 1. We can allocate these objects among agents 2 and 3 in any way we like with a single shared object, and use the other objects (which are less valuable) to compensate agent 1 without additional splits (see Case 0 in Appendix 0.A);
  • $U_{4}<U_{1}\leq\frac{2}{3}\cdot V$ — there is always a PROP allocation. The intuitive meaning of this case is that there is a set of objects objects (namely ${\cal X}_{1}$) that are worth a lot to all three agents, but the sum of these objects is at most $2/3$ of the total value $V$. Therefore, it is enough to allocate this set among two of the three agents (giving each of them at least $V/3$), and use the other objects to compensate the third agent (see Case 0 in Appendix 0.A);
  • $U_{1}>\frac{2}{3}\cdot V$. The intuitive meaning of this case is that there is a set of objects objects (namely ${\cal X}_{1}$) that are worth a lot to all three agents, and their sum is larger than $2/3$ the total value. Therefore, we must partition these objects among all three agents. To do this, we use MaxMinIdentical$(3,1)$. If the largest bundle sum is at most $\frac{U_{1}+U_{2}+2\cdot(U_{3}+U_{4}+U_{5}+U_{6})}{3}$, then a PROP allocation exists and we answer “yes”; otherwise, we prove that no PROP allocation can exist, so we answer “no” (see Case 0 in Appendix 0.A).
This is a polynomial time algorithm for Prop-SharedObj$(3,1)$ with equal-sum generalized binary utilities, proving Theorem 4.1.
Open problem 2.
Given some fixed $n\geq 3$, what is the run-time complexity of the problems Prop-SharedObj$(n,n-2)$ and of EF-SharedObj$(n,n-2)$ for generalized binary and arbitrary utilities?
In particular, what is the run-time complexity of Prop-SharedObj$(3,1)$ for arbitrary utilities and EF-SharedObj$(3,1)$ for generalized binary utilities?
The same questions are interesting for CONS and EQ.
5 Non-Degenerate Valuations
NOTE: following reviewers’ comments, this section was rewritten in a substantially more formal way. Most of the proofs have been moved to appendices and can be found in the full version [20].
As mentioned in the introduction, Sandomirskiy and Segal-Halevi [82] prove that problems PropFpo-Sharings$(n,s)$, EFFpo-Sharings$(n,s)$, PropFpo-SharedObj$(\allowbreak n,s)$, EFFpo-SharedObj$(n,s)$ are solvable in time $O(poly(m))$ for any fixed $n$ and $s$, whenever the valuations are non-degenerate, and argue — somewhat informally — that “almost all” valuations are non-degenerate; this means that almost all instances of the above problems are easy. In this section we would like to show that the requirement of fractional Pareto-optimality (fPO) is essential for this result: when this requirement is dropped, or even just relaxed to dPO, computational hardness strikes even for non-degenerate valuations, and it is no longer true that almost all instances are easy.
To prove this statement, we first have to formally define the notion of “almost all valuations are easy”.
5.1 Definitions
We consider a decision problem $P$, whose input is a vector of some $t$ non-negative integers. As the input size is measured by $t$, the binary encoding length of the numbers should be polynomial in $t$. For simplicity, we assume that all input integers are in the range $[0,2^{ct}]$, for some constant $c$ that may depend on the problem. We denote the set of possible inputs of size $t$ (that is, $t$-sized vectors of integers in $[0,2^{ct}]$) by $\mathcal{I}({t,c})$.
For any input vector $x\in\mathcal{I}({t,c})$, we denote its infinity norm by $\|x\|$. For any $x$ and integer $r$, we denote by $\mathcal{B}(x,r)$ the ball of radius $r$ around $x$: $\mathcal{B}(x,r):=\{x^{\prime}\in\mathcal{I}({t,c}):\|x-x^{\prime}\|\leq r\}$.
The notion of “almost all instances are easy” is formalized by the following definition of “generically polynomial-time algorithm”.
Definition 2.
Given an algorithm $A$ for a problem $P$, we say that $A$ runs generically in polynomial time if there exists a polynomial function $f_{p}$ such that, for every size $t$ and input $x\in\mathcal{I}({t,c})$, there is a subset of “Good inputs” $G(x)\subseteq\mathcal{B}(x,f_{p}(t))$, such that the following holds:
(a) Algorithm $A$ runs in time poly($t$) on all inputs in $G(x)$;
(b) The fraction of good inputs approaches 1, that is,
$\lim_{t\to\infty}\min_{x\in\mathcal{I}({t,c})}\frac{|G(x)|}{|\mathcal{B}(x,f_{p}(t))|}=1,$
(c) Given $x$, it is possible to compute in time poly($t$) a vector in $G(x)$.
The results in [82] imply:
Proposition 1.
For every fixed $n,s$, decision problems PropFpo-Sharings$(n,s)$,EFFpo-Sharings$(n,s)$,PropFpo-SharedObj$(n,s)$,EFFpo-SharedObj$(n,s)$ have algorithms that run in generically-polynomial time.
Proof.
In these problems, the input is $x\equiv v=$ a valuation matrix, and the number of integers in the input is $t=mn=$ the number of values in the valuation matrix. We choose the polynomial $f_{p}(t):=t^{3}$. We define the set $G(v)$ as the subset of matrices in $\mathcal{B}(v,t^{3})$ that define non-degenerate valutions. We show that this set satisfies Definition 2.
(a) Is satisfied by the algorithms in [82].
[…]
$\frac{|G(v)|}{|\mathcal{B}(v,t^{3})|}$ $\geq\frac{\prod_{i,o}(r_{i,o}-t)}{\prod_{i,o}(r_{i,o})}$
$={\prod_{i,o}(1-t/r_{i,o})}\geq(1-t/(t^{3}+1))^{t},$
which approaches $1$ as $t\to\infty$; this formalizes the claim that “almost all inputs are non-degenerate”.
(c) To compute an input in $G(v)$, it is sufficient to check at most $t$ options for changing each coefficient $v_{i,o}$; this can be done using polynomial in $t$ time. ∎
Our goal is to prove that some problems do not have generically-polynomial-time algorithms. To this end, we first define a multi-reduction. We present a simplified version first, and then the full version.
Definition 3 (multi-reduction – simplified).
Given two decision problems $P_{1}$ and $P_{2}$, both defined on inputs in $\mathcal{I}({t,c})$ for some $c\geq 1$, a polynomial-time multi-reduction from $P_{1}$ to $P_{2}$ is a family of functions, $h_{t}:\mathcal{I}({t,c})\to\mathcal{I}({t,c})$, which maps an input for $P_{1}$ to an input for $P_{2}$, and satisfies the following:
(a) $h_{t}$ runs in time poly$(t)$;
(b) There exists a super-polynomial function $f_{e}$ such that, for all $t$ and all $x_{1}\in\mathcal{I}({t,c})$, when $x_{2}:=h_{t}(x_{1})$,
[…]
A multi-reduction is stronger than a usual reduction in that each input to $P_{1}$ is transformed simultaneously to an super-polynomially-large set of inputs to $P_{2}$, all of which have the same output.
Definition 3 is “simplified” since it assumes that the size and encoding length of the inputs to $P_{1}$ is equal to the size and encoding length of the inputs to $P_{2}$. In fact, reductions often add some inputs or increase the encoding length. We handle this technical issue by assuming that the input $x_{1}$ is in $\mathcal{I}({t_{1},c_{1}})$ and the transformed input $x_{2}$ is in $\mathcal{I}({t_{2},c_{2}})$, where $t_{2}$ depends polynomially on $t_{1}$ but they do not have to be equal. We also allow the constants $c_{1}$ and $c_{2}$ to differ.
Definition 4 (multi-reduction – full).
Given two decision problems, $P_{1}$ defined on inputs in $\mathcal{I}({t_{1},c_{1}})$ for some constant $c_{1}\geq 1$ and $P_{2}$ defined on inputs in $\mathcal{I}({t_{2},c_{2}})$ for some constant $c_{2}\geq 1$, a multi-reduction from $P_{1}$ to $P_{2}$ consists of a polynomial-time-computable function $t_{2}:\mathbb{N}\to\mathbb{N}$ and a family of functions, $h_{t_{1}}:\mathcal{I}({t_{1},c_{1}})\to\mathcal{I}({t_{2}(t_{1}),c_{2}})$, which maps an input for $P_{1}$ to an input for $P_{2}$, and satisfies the following:
(a) $h_{t_{1}}$ runs in time poly$(t_{1})$;
(b) There exists a super-polynomial function $f_{e}$ such that, for all $t_{1}$ and all $x_{1}\in\mathcal{I}({t_{1},c_{1}})$, when $x_{2}:=h_{t_{1}}(x_{1})$,
[…]
Definition 5.
A decision problem $P_{2}$ is called generically NP hard if there exists a multi-reduction from some NP-hard problem $P_{1}$ to $P_{2}$.
In Appendix 0.E we prove that the relation between “generically-polynomial” and “generically-NP-hard” is analogous to the relation between “polynomial” and “NP-hard”:
Proposition 2.
If a problem is generically-NP-hard, then it does not have a generically-polynomial algorithm unless P=NP.
Now we are ready for our main results: showing that fair division problems without the fPO requirement are generically-NP-hard, and therefore probably do not have generically-polynomial-time algorithms (in particular, they are NP-hard even for non-degenerate valuations).
5.2 EF and PROP allocations — fixed $n$
Theorem 5.1.
The decision problems Prop-Sharings$(n,s)$ and EF-Sharings$(n,s)$ are generically NP-hard in the following cases:
(a) For any fixed $n\geq 2$ and $s=0$;
(b) For any fixed $n\geq 3$ and $s\in\{0,\dots,n-3\}$.
The proof is by showing a multi-reduction from $k$-way partition, where $k=n$ in part (a) and $k=n-s-1$ in part (b). Due to space constraints, we moved all proofs to Appendix 0.E.
We could not adapt the proof of Theorem 5.1 to shared objects. We can still prove similar results for shared objects, but with a different multi-reduction.
Theorem 5.2.
Problems Prop-SharedObj$(n,s)$ and EF-SharedObj$(n,s)$ are generically NP-hard in the following cases:
(a) For any fixed $n\geq 2$ and $s=0$;
(b) For any fixed $n\geq 3$ and $s\in\{0,\dots,n-3\}$.
The hardness of Theorem 5.1(a) and Theorem 5.2(a) remains even if we add the requirement of discrete PO (in contrast to fractional PO). We consider only part (a) as the requirement of discrete PO makes sense only for allocations without sharing.
Theorem 5.3.
For any fixed integer $n\geq 2$, the decision problems PropDpo-Sharings$(n,0)$ ($\equiv$ PropDpo-SharedObj$(n,0)$) and EFDpo-Sharings$(n,0)$ ($\equiv$ EFDpo-SharedObj$(n,0)$) are generically NP-hard.
Theorem 5.3 is interesting as it shows a crucial difference between the apparently-similar concepts fPO and dPO: whilst the fPO allocations can be enumerated in polynomial time (as their number is polynomial in $m$ when the valuations are non-degenerate [82]), the dPO allocations cannot.
When $s=n-1$, an EF and PROP allocation with $s$ sharings always exists (see Appendix 0.B). Therefore, only the case $s=n-2$ remains open. With identical valuations, this case is NP-hard with sharings and polynomial with shared objects [21]. We do not know if the same is true with non-degenerate valuations.
Open problem 3.
For any $n\geq 3$ and $s=n-2$, do the problems Prop-Sharings$(n,s)$, EF-Sharings$(n,s)$, Prop-SharedObj$(n,s)$, EF-SharedObj$(n,s)$ have a generically-polynomial-time algorithm?
For equitability and consensus allocation, we do not yet have any generic results:
Open problem 4.
(a) For any $n\geq 2$ and $s\in[0,n-1)$, do the problems Eq-Sharings$(n,s)$, Eq-SharedObj$(n,s)$ have a generically-polynomial-time algorithm?
(b) For any $n\geq 2$ and $s\in[0,n(n-1))$, do the problems Cons-Sharings$(n,s)$, Cons-SharedObj$(n,s)$ have a generically-polynomial-time algorithm?
5.3 EF and PROP allocations — unbounded $n$
So far we assumed that $n$ is fixed. If $n$ is unbounded (part of the input), then we can prove strong NP-hardness for both sharings and shared objects. This requires to adapt the notions of multi-reduction and generic NP-hardness to strong NP-hardness. All definitions and proofs are in Appendix 0.F.
6 Conclusion
Our work presented several results related to fairly allocating objects among agents, with a mixture of divisible and indivisible objects. In our work all objects are divisible, however, divisibility is highly discouraged. We covered many fairness and efficiency concepts, with a bound on the number of sharings or shared objects. We tackle the difficulty of finding an allocation for agents with arbitrary valuations by restricting the agent valuations to some well-studied domains, especially the binary, generalized-binary, and the non-degenerate valuations. In addition, our work shows that sometimes a polynomial algorithm can be designed if we allow a bounded number of objects to be shared among agents. Such a behavior is already analyzed in [21], and our work can be seen as a confirmation — when searching for a fair allocation is too hard, considering the objects as divisible, but still constraining the number of shared objects, might be reasonable for the agents, and may significantly decrease the runtime of the search. Thus, we highlighted a new way to relax the difficult fair-division problem, as it is common in the literature using for example Envy-freeness up to one good (EF1).
Our main result considers 3 agents under equal-sum generalized binary utilities. Despite the fact that we consider only 3 agents, the algorithm is complicated, and involves an exhaustive cases analysis. That is why, as a next challenge, one can try to generalize our algorithm and see if it is possible to extend it for any fixed number $n$ of agents. Also, it is interesting to see if the same result holds for agents under general additive utilities.
We dedicated one section to non-degenerate valuations. Despite the surprising polynomial time algorithm designed in [82], several results in our paper point up the hardness of this problem, even when allowing shared objects.
Along the paper, we left many open problems. This paper is an invitation for researchers interested in the fair-division field, to find new results, with the goal to find theoretic and practical solutions, to numerous fair-division problems.
Acknowledgements.
This work was inspired by Achikam Bitan, Steven Brams and Shahar Dobzinski, who expressed their dissatisfaction with the current trend of approximate-fairness (SCADA conference, Weizmann Institute, Israel 2018). We are grateful to participants of De Aequa Divisione Workshop on Fair Division (LUISS, Rome, 2019), Workshop on Theoretical Aspects of Fairness (WTAF, Patras, 2019) and the rationality center game theory seminar (HUJI, Jerusalem, 2019) for their helpful comments. Suggestions of Herve Moulin, Antonio Nicolo, Nisarg Shah and Rohit Vaish were especially useful.
Several members of the stack-exchange network provided helpful answers, in particular: D.W, Peter Taylor, Gamow, Sasho Nikolov, Chao Xu, Mikhail Rudoy, xskxzr, Philipp Christophel, Kevin Dalmeijer, Bodo Manthey, and Niklas Rieken.
We are grateful to the anonymous referees of WTAF 2019 and SAGT 2024 for their helpful comments.
This work was partly funded by Israel Science Foundation grant no. 712/20.
APPENDIX
Appendix 0.A Generalized Binary Utilities – Cases Details
In this section, we handle the 3 cases to prove Theorem 4.1. We consider three cases:
  • $U_{1}\leq U_{4}$ — there is always a PROP allocation (Case 0);
  • $U_{4}<U_{1}\leq\frac{2}{3}\cdot V$ — there is always a PROP allocation (Case 0);
  • $U_{1}>\frac{2}{3}\cdot V$ — there is not always a PROP allocation (Case 0).
Case 0 $U_{1}\leq U_{4}$, there is always a PROP allocation
${\cal X}_{j}$ $\sum_{x_{i}\in{\cal X}_{j}}u_{1}(x_{i})$ $\sum_{x_{i}\in{\cal X}_{j}}u_{2}(x_{i})$ $\sum_{x_{i}\in{\cal X}_{j}}u_{3}(x_{i})$
${\cal X}_{1}$ $U_{1}$ $U_{1}$ $U_{1}*$
${\cal X}_{2}$ $U_{2}$ $U_{2}*$ 0
${\cal X}_{3}$ $U_{3}*$ 0 $U_{3}$
${\cal X}_{4}$ 0 $U_{4}\longrightarrow\frac{U_{1}+U_{4}}{3}*$ $U_{4}\longrightarrow\frac{2\cdot U_{4}-U_{1}}{3}*$
${\cal X}_{5}$ $U_{5}*$ 0 0
${\cal X}_{6}$ 0 $U_{6}*$ 0
${\cal X}_{7}$ 0 0 $U_{7}*$
bundle sum $U_{3}+U_{5}$ $U_{2}+\frac{U_{1}+U_{4}}{3}+U_{6}$ $U_{1}+\frac{2\cdot U_{4}-U_{1}}{3}+U_{7}$
We assign the objects as shown in the table, such that there is at most one shared object in ${\cal X}_{4}$. The intuitive meaning of this case is that a set of objects is worth a lot to agents 2 and 3, but worth little to agent 1. We can allocate these objects among agents 2 and 3 in any way we like with a single shared object, and use the other objects (which are less valuable) to compensate agent 1 without additional splits. Recall that by assumption $U_{2}\leq U_{3}\leq U_{4}$ and $U_{1}\leq U_{4}$. Formally, the allocation is PROP since,
  • $U_{3}+U_{5}=U_{4}+U_{6}\geq\frac{U_{4}+U_{4}+U_{4}+U_{6}}{3}\geq\frac{U_{1}+U_{2}+U_{4}+U_{6}}{3}=\frac{V}{3}$;
  • $U_{2}+\frac{U_{1}+U_{4}}{3}+U_{6}\geq\frac{U_{1}+U_{2}+U_{4}+U_{6}}{3}=\frac{V}{3}$;
  • $U_{1}+\frac{2\cdot U_{4}-U_{1}}{3}+U_{7}\geq\frac{U_{1}+U_{4}+U_{4}+U_{7}}{3}\geq\frac{U_{1}+U_{3}+U_{4}+U_{7}}{3}=\frac{V}{3}$.
Case 0 $U_{4}<U_{1}\leq\frac{2}{3}\cdot V$, there is always a PROP allocation
${\cal X}_{j}$ $\sum_{x_{i}\in{\cal X}_{j}}u_{1}(x_{i})$ $\sum_{x_{i}\in{\cal X}_{j}}u_{2}(x_{i})$ $\sum_{x_{i}\in{\cal X}_{j}}u_{3}(x_{i})$
${\cal X}_{1}$ $U_{1}$ $U_{1}\longrightarrow U_{1}^{\prime}*$ $U_{1}\longrightarrow U_{1}^{\prime\prime}*$
${\cal X}_{2}$ $U_{2}*$ $U_{2}$ 0
${\cal X}_{3}$ $U_{3}*$ 0 $U_{3}$
${\cal X}_{4}$ 0 $U_{4}\longrightarrow U_{4}^{\prime}*$ $U_{4}\longrightarrow U_{4}^{\prime\prime}*$
${\cal X}_{5}$ $U_{5}*$ 0 0
${\cal X}_{6}$ 0 $U_{6}*$ 0
${\cal X}_{7}$ 0 0 $U_{7}*$
bundle sum $U_{2}+U_{3}+U_{5}$ $U_{1}^{\prime}+U_{4}^{\prime}+U_{6}$ $U_{1}^{\prime\prime}+U_{4}^{\prime\prime}+U_{7}$
We assign the objects as shown in the table, such that there is at most one shared object in ${\cal X}_{1}\cup{\cal X}_{4}$. We combine ${\cal X}_{1}$ and ${\cal X}_{4}$, and cut at half such that $U_{1}^{\prime}+U_{4}^{\prime}=U_{1}^{\prime\prime}+U_{4}^{\prime\prime}=\frac{U_{1}+U_{4}}{2}$. The intuitive meaning of this case is that there is a set of objects that are worth a lot to all three agents, but the sum of these objects is at most $2/3$ of the total value $V$. Therefore, it is enough to allocate this set among two of the three agents (giving each of them at least $V/3$), and use the other objects to compensate the third agent. Recall that by assumption $U_{2}\leq U_{3}\leq U_{4}<U_{1}$, and, $U_{1}\leq\frac{2}{3}\cdot V$. Formally, the allocation is PROP since,
  • $U_{2}+U_{3}+U_{5}=U_{1}+U_{2}+U_{3}+U_{5}-U_{1}=V-U_{1}\geq V-\frac{2}{3}\cdot V=\frac{V}{3}$;
  • $U_{1}^{\prime}+U_{4}^{\prime}+U_{6}=\frac{U_{1}+U_{4}}{2}+U_{6}=\frac{U_{1}+U_{4}}{3}+\frac{U_{1}+U_{4}}{6}+U_{6}\geq\frac{U_{1}+U_{4}+U_{6}}{3}+\frac{U_{2}+U_{2}}{6}\geq\frac{U_{1}+U_{2}+U_{4}+U_{6}}{3}=\frac{V}{3}$;
  • $U_{1}^{\prime\prime}+U_{4}^{\prime\prime}+U_{7}=\frac{U_{1}+U_{4}}{2}+U_{7}=\frac{U_{1}+U_{4}}{3}+\frac{U_{1}+U_{4}}{6}+U_{7}\geq\frac{U_{1}+U_{4}+U_{7}}{3}+\frac{U_{3}+U_{3}}{6}\geq\frac{U_{1}+U_{3}+U_{4}+U_{7}}{3}=\frac{V}{3}$.
Case 0 $U_{1}>\frac{2}{3}\cdot V$
The intuitive meaning of this case is that there is a set of objects that are worth a lot to all three agents, and their sum is larger than $2/3$ the total value. Therefore, we must partition these objects (objects in ${\cal X}_{1}$) among all three agents, using MaxMinIdentical$(3,1)$. We run MaxMinIdentical$(3,1)$ on objects in ${\cal X}_{1}$, and we denote the resulting bundles sum by $b_{1},b_{2},b_{3}$, where without loss of generality, $b_{3}$ is the largest bundle sum. By Lemma 1, $b_{1}=b_{2}\leq b_{3}$.
We consider three cases,
  • $b_{3}\leq\frac{U_{1}+U_{3}+U_{4}-2\cdot U_{7}}{3}$;
  • $\frac{U_{1}+U_{3}+U_{4}-2\cdot U_{7}}{3}<b_{3}\leq\frac{U_{1}+U_{2}+2\cdot(U_{3}+U_{4}+U_{5}+U_{6})}{3}$;
  • $b_{3}>\frac{U_{1}+U_{2}+2\cdot(U_{3}+U_{4}+U_{5}+U_{6})}{3}$.
Case 0.1   $b_{3}\leq\frac{U_{1}+U_{3}+U_{4}-2\cdot U_{7}}{3}$, there is always a PROP allocation
${\cal X}_{j}$ $\sum_{x_{i}\in{\cal X}_{j}}u_{1}(x_{i})$ $\sum_{x_{i}\in{\cal X}_{j}}u_{2}(x_{i})$ $\sum_{x_{i}\in{\cal X}_{j}}u_{3}(x_{i})$
${\cal X}_{1}$ $U_{1}\longrightarrow U_{1}^{\prime}*$ $U_{1}\longrightarrow U_{1}^{\prime\prime}*$ $U_{1}\longrightarrow U_{1}^{\prime\prime\prime}*$
${\cal X}_{2}$ $U_{2}$ $U_{2}*$ 0
${\cal X}_{3}$ $U_{3}*$ 0 $U_{3}$
${\cal X}_{4}$ 0 $U_{4}$ $U_{4}*$
${\cal X}_{5}$ $U_{5}*$ 0 0
${\cal X}_{6}$ 0 $U_{6}*$ 0
${\cal X}_{7}$ 0 0 $U_{7}*$
bundle sum $U_{1}^{\prime}+U_{3}+U_{5}$ $U_{1}^{\prime\prime}+U_{2}+U_{6}$ $U_{1}^{\prime\prime\prime}+U_{4}+U_{7}$
We assign the objects as shown in the table. The intuitive meaning of this is case is that the partition of $U_{1}$ is quite balanced, so we can make it PROP by allocating the other objects.
Denote the value of the most valuable object in $U_{1}$ by $3M$. $b_{1}=b_{2}=b_{3}$ and $U_{4}<3M$
By Lemma 2, the shared object is the most valuable one (with value $3M$), by the pigeonhole principle, there is at least one bundle in which the part of the shared object is at least $M$. Give this bundle to agent 3.
We have $\frac{U_{4}}{3}<M$, so, starting from the perfect allocation, we can move $\frac{U_{4}}{3}$ from the shared object from agent 3 to agent 2. Therefore, $U_{1}^{\prime}=\frac{U_{1}}{3}$, $U_{1}^{\prime\prime}=\frac{U_{1}+U_{4}}{3}$ and $U_{1}^{\prime\prime\prime}=\frac{U_{1}-U_{4}}{3}$.
The allocation is PROP since,
  • $U_{1}^{\prime}+U_{3}+U_{5}=\frac{U_{1}}{3}+U_{3}+U_{5}\geq\frac{U_{1}+U_{2}+U_{3}+U_{5}}{3}=\frac{V}{3}$;
  • $U_{1}^{\prime\prime}+U_{2}+U_{6}=\frac{U_{1}+U_{4}}{3}+U_{2}+U_{6}\geq\frac{U_{1}+U_{2}+U_{4}+U_{6}}{3}=\frac{V}{3}$;
  • $U_{1}^{\prime\prime\prime}+U_{4}+U_{7}=\frac{U_{1}-U_{4}}{3}+U_{4}+U_{7}\geq\frac{U_{1}+U_{3}+U_{4}+U_{7}}{3}=\frac{V}{3}$.
$b_{1}=b_{2}=b_{3}$ and $U_{4}\geq 3M$
Since the largest object in $U_{1}$ is equal to $3M$, we define $U_{1}^{\prime\prime\prime}$ as follows. Cut from ${\cal X}_{1}$ a subset with value $\frac{U_{1}+U_{3}-2\cdot(U_{4}+U_{7})}{3}$. The cut is feasible since:
$U_{1}$ $>\frac{2}{3}\cdot V\iff 3\cdot U_{1}>2\cdot V=2\cdot(U_{1}+U_{3}+U_{4}+U_{7})\iff U_{1}>2\cdot(U_{3}+U_{4}+U_{7});$ (1)
so $0<\frac{U_{1}+U_{3}-2\cdot(U_{4}+U_{7})}{3}\leq U_{1}$. We add the shared object (if any) to this subset. Now there is no shared object in $U_{1}^{\prime\prime\prime}$, and
$\frac{U_{1}+U_{3}-2\cdot(U_{4}+U_{7})}{3}\leq U_{1}^{\prime\prime\prime}\leq\frac{U_{1}+U_{3}-2\cdot(U_{4}+U_{7})}{3}+3M\leq\frac{U_{1}+U_{3}-2\cdot(U_{4}+U_{7})}{3}+U_{4}$
$=\frac{U_{1}+U_{3}+U_{4}-2\cdot U_{7}}{3}.$
The sum of the remaining objects is at least:
$U_{1}-U_{1}^{\prime\prime\prime}\geq U_{1}-\frac{U_{1}+U_{3}+U_{4}-2\cdot U_{7}}{3}=\frac{2\cdot(U_{1}+U_{7})-U_{3}-U_{4}}{3}.$
Since there is no shared object in $U_{1}^{\prime\prime\prime}$ we can cut the remaining objects and set $U_{1}^{\prime\prime}=\frac{U_{1}+U_{4}-2\cdot(U_{3}+U_{7})}{3}$. The cut is feasible since:
[…]
$\geq\frac{U_{1}+U_{4}-2\cdot(U_{3}+U_{7})}{3};$
so $0<\frac{U_{1}+U_{4}-2\cdot(U_{3}+U_{7})}{3}\leq U_{1}-U_{1}^{\prime\prime\prime}$. Also, $U_{1}^{\prime}\geq U_{1}-U_{1}^{\prime\prime\prime}-U_{1}^{\prime\prime}$, which is at least:
$U_{1}^{\prime}\geq\frac{2\cdot(U_{1}+U_{7})-U_{3}-U_{4}}{3}-\frac{U_{1}+U_{4}-2\cdot(U_{3}+U_{7})}{3}=\frac{U_{1}+U_{3}-2\cdot U_{4}+4\cdot U_{7}}{3}.$
The allocation is PROP since,
  • $U_{1}^{\prime}+U_{3}+U_{5}\geq\frac{U_{1}+U_{3}-2\cdot U_{4}+4\cdot U_{7}}{3}+U_{3}+U_{5}\stackrel{{\scriptstyle(U_{3}+U_{5}=U_{4}+U_{6})}}{{\geq}}\frac{U_{1}+U_{3}-2\cdot U_{4}+U_{7}}{3}+U_{4}\geq\frac{U_{1}+U_{3}+U_{4}+U_{7}}{3}=\frac{V}{3}$;
  • $U_{1}^{\prime\prime}+U_{2}+U_{6}\geq\frac{U_{1}+U_{4}-2\cdot(U_{2}+U_{6})}{3}+U_{2}+U_{6}\geq\frac{U_{1}+U_{2}+U_{4}+U_{6}}{3}=\frac{V}{3}$;
  • $U_{1}^{\prime\prime\prime}+U_{4}+U_{7}\geq\frac{U_{1}+U_{3}-2\cdot(U_{4}+U_{7})}{3}+U_{4}+U_{7}\geq\frac{U_{1}+U_{3}+U_{4}+U_{7}}{3}=\frac{V}{3}$.
$b_{1}=b_{2}<b_{3}$
Set $U_{1}^{\prime\prime\prime}=b_{3}$. By assumption, $\frac{U_{1}}{3}<b_{3}=U_{1}^{\prime\prime\prime}\leq\frac{U_{1}+U_{3}+U_{4}-2\cdot U_{7}}{3}$, and by Lemma 1, there is no shared object in $U_{1}^{\prime\prime\prime}$. So, we reach the exact same structure from Case 0.1. Using the same arguments the allocation is PROP.
Case 0.2   $\frac{U_{1}+U_{3}+U_{4}-2\cdot U_{7}}{3}\leq b_{3}\leq\frac{U_{1}+U_{2}+2\cdot(U_{3}+U_{4}+U_{5}+U_{6})}{3}$, there is always a PROP allocation
${\cal X}_{j}$ $\sum_{x_{i}\in{\cal X}_{j}}u_{1}(x_{i})$ $\sum_{x_{i}\in{\cal X}_{j}}u_{2}(x_{i})$ $\sum_{x_{i}\in{\cal X}_{j}}u_{3}(x_{i})$
${\cal X}_{1}$ $U_{1}\longrightarrow U_{1}^{\prime}*$ $U_{1}\longrightarrow U_{1}^{\prime\prime}*$ $U_{1}\longrightarrow U_{1}^{\prime\prime\prime}*$
${\cal X}_{2}$ $U_{2}\longrightarrow U_{2}^{\prime}*$ $U_{2}\longrightarrow U_{2}^{\prime\prime}*$ 0
${\cal X}_{3}$ $U_{3}*$ 0 $U_{3}$
${\cal X}_{4}$ 0 $U_{4}*$ $U_{4}$
${\cal X}_{5}$ $U_{5}*$ 0 0
${\cal X}_{6}$ 0 $U_{6}*$ 0
${\cal X}_{7}$ 0 0 $U_{7}*$
bundle sum $U_{1}^{\prime}+U_{2}^{\prime}+U_{3}+U_{5}$ $U_{1}^{\prime\prime}+U_{2}^{\prime\prime}+U_{4}+U_{6}$ $U_{1}^{\prime\prime\prime}+U_{7}$
We assign the objects as shown in the table. This case is an intermediate case, where the partition of $U_{1}$ is not balanced, but the maximum is still not very high. Then, we can satisfy one of the agents by giving him the largest part of $U_{1}$, and compensate the other agents using the remaining objects.
Set $U_{1}^{\prime\prime\prime}=b_{3}$. By assumption, $b_{3}=U_{1}^{\prime\prime\prime}\leq\frac{U_{1}+U_{2}+2\cdot(U_{3}+U_{4}+U_{5}+U_{6})}{3}$, so the sum of the remaining objects is at least:
$U_{1}^{\prime}+U_{1}^{\prime\prime}+U_{2}^{\prime}+U_{2}^{\prime\prime}=U_{1}-U_{1}^{\prime\prime\prime}+U_{2}$ $\geq U_{1}-\frac{U_{1}+U_{2}+2\cdot(U_{3}+U_{4}+U_{5}+U_{6})}{3}+U_{2}$
$=\frac{2\cdot(U_{1}+U_{2})-2\cdot(U_{3}+U_{4}+U_{5}+U_{6})}{3}.$
By Lemma 1, there is no shared object in $U_{1}^{\prime\prime\prime}$, so we can cut the remaining objects and set $U_{1}^{\prime}+U_{2}^{\prime}=\frac{U_{1}+U_{2}-2\cdot(U_{3}+U_{5})}{3}$, and the remainder $U_{1}^{\prime\prime}+U_{2}^{\prime\prime}\geq\frac{U_{1}+U_{2}-2\cdot(U_{4}+U_{6})}{3}$. The allocation is PROP since,
  • $U_{1}^{\prime}+U_{2}^{\prime}+U_{3}+U_{5}\geq\frac{U_{1}+U_{2}-2\cdot(U_{3}+U_{5})}{3}+U_{3}+U_{5}=\frac{U_{1}+U_{2}+U_{3}+U_{5}}{3}=\frac{V}{3}$;
  • $U_{1}^{\prime\prime}+U_{2}^{\prime\prime}+U_{4}+U_{6}\geq\frac{U_{1}+U_{2}-2\cdot(U_{4}+U_{6})}{3}+U_{4}+U_{6}=\frac{U_{1}+U_{2}+U_{4}+U_{6}}{3}=\frac{V}{3}$;
  • $U_{1}^{\prime\prime\prime}+U_{7}\geq\frac{U_{1}+U_{3}+U_{4}-2\cdot U_{7}}{3}+U_{7}=\frac{U_{1}+U_{3}+U_{4}+U_{7}}{3}=\frac{V}{3}$.
Case 0.3   $b_{3}>\frac{U_{1}+U_{2}+2\cdot(U_{3}+U_{4}+U_{5}+U_{6})}{3}$, there is no PROP allocation
In this case, the optimal partition of $U_{1}$ is so unbalanced, that it is not possible to make it PROP using the remaining objects.
Lemma 3.
If there exists a PROP allocation, then the algorithm for MaxMinIdentical$(3,1)$ must find a partition with
$b_{3}\leq\frac{U_{1}+U_{2}+2\cdot(U_{3}+U_{4}+U_{5}+U_{6})}{3}$.
Proof.
Suppose there exists a PROP allocation. For each agent $i$ and subset ${\cal X}_{j}$, denote by $F_{i}(U_{j})$ the total value of objects assigned to $i$ from subset ${\cal X}_{j}$ in that allocation. We now show that, for all $i\in\{1,2,3\}$, $F_{i}(U_{1})\leq\frac{U_{1}+U_{2}+2\cdot(U_{3}+U_{4}+U_{5}+U_{6})}{3}$.
First assume that $F_{3}(U_{1})\geq F_{2}(U_{1}),F_{1}(U_{1})$. By PROP definition, we know that:
$F_{1}(U_{1})+F_{1}(U_{2})+F_{1}(U_{3})+U_{5}\geq\frac{U_{1}+U_{2}+U_{3}+U_{5}}{3};$
$F_{2}(U_{1})+F_{2}(U_{2})+F_{2}(U_{4})+U_{6}\geq\frac{U_{1}+U_{2}+U_{4}+U_{6}}{3}.$
So,
$F_{1}(U_{1})\geq\frac{U_{1}+U_{2}+U_{3}-2\cdot U_{5}}{3}-F_{1}(U_{2})-F_{1}(U_{3});$
$F_{2}(U_{1})\geq\frac{U_{1}+U_{2}+U_{4}-2\cdot U_{6}}{3}-F_{2}(U_{2})-F_{2}(U_{4}).$
Summing up these inequalities gives:
$F_{1}(U_{1})+F_{2}(U_{1})$ $\geq\frac{U_{1}+U_{2}+U_{3}-2\cdot U_{5}}{3}-F_{1}(U_{2})-F_{1}(U_{3})+\frac{U_{1}+U_{2}+U_{4}-2\cdot U_{6}}{3}-F_{2}(U_{2})-F_{2}(U_{4})$
$\geq\frac{2\cdot(U_{1}+U_{2})+U_{3}+U_{4}-2\cdot(U_{5}+U_{6})}{3}-U_{2}-U_{3}-U_{4}$
$=\frac{2\cdot U_{1}-U_{2}-2\cdot(U_{3}+U_{4}+U_{5}+U_{6})}{3}.$
Since $F_{1}(U_{1})+F_{2}(U_{1})+F_{3}(U_{1})=U_{1}$, we have $F_{1}(U_{1})+F_{2}(U_{1})=U_{1}-F_{3}(U_{1})$, so:
$U_{1}-F_{3}(U_{1})$ $\geq\frac{2\cdot U_{1}-U_{2}-2\cdot(U_{3}+U_{4}+U_{5}+U_{6})}{3}$
$F_{3}(U_{1})$ $\leq U_{1}-\frac{2\cdot U_{1}-U_{2}-2\cdot(U_{3}+U_{4}+U_{5}+U_{6})}{3}$
$\leq\frac{U_{1}+U_{2}+2\cdot(U_{3}+U_{4}+U_{5}+U_{6})}{3}.$
Using a similar logic, if $F_{2}(U_{1})\geq F_{3}(U_{1}),F_{1}(U_{1})$, we have
$F_{2}(U_{1})$ $\leq\frac{U_{1}+2\cdot U_{2}+U_{3}+2\cdot U_{4}+2\cdot U_{5}+2\cdot U_{7}}{3}$
$=\frac{U_{1}+U_{2}+2\cdot(U_{3}+U_{4}+U_{5}+U_{6})}{3}+\frac{U_{2}+2\cdot U_{7}-U_{3}-2\cdot U_{6}}{3}.$
Since $U_{2}\leq U_{3}$ and $U_{7}\leq U_{6}$, $F_{2}(U_{1})\leq\frac{U_{1}+U_{2}+2\cdot(U_{3}+U_{4}+U_{5}+U_{6})}{3}$.
Again, using similar arguments, if $F_{1}(U_{1})\geq F_{2}(U_{1}),F_{3}(U_{1})$, we have
$F_{1}(U_{1})$ $\leq\frac{U_{1}+2\cdot U_{2}+2\cdot U_{3}+U_{4}+2\cdot U_{6}+2\cdot U_{7}}{3}$
$=\frac{U_{1}+U_{2}+2\cdot(U_{3}+U_{4}+U_{5}+U_{6})}{3}+\frac{U_{2}+U_{7}-U_{4}-2\cdot U_{5}}{3}.$
Since $U_{2}\leq U_{4}$ and $U_{7}\leq U_{5}$, $F_{1}(U_{1})\leq\frac{U_{1}+U_{2}+2\cdot(U_{3}+U_{4}+U_{5}+U_{6})}{3}$. ∎
Appendix 0.B Worst-case Upper Bounds on Sharing
Several previous works proved worst-case upper bounds were proved using different combinatorial techniques. For completeness, we present here alternative, simpler proofs for all these results using a more general technique: linear programming.
0.B.1 Existence of upper bounds
In this section, we show that for every fairness criterion F among our four criteria, there is a function $S_{F}(n)$ such that any instance with $n$ agents has an “F” allocation with at most $S_{F}(n)$ sharings. Most of the results below were proved previously using combinatorial techniques [25, 14, 82]. We present below alternative proofs using a more general technique. The main tool is the following known fact: a bounded feasible linear program in equational form, with $k$ equational constraints, has a basic feasible solution — a solution in which at most $k$ variables are non-zero [72]. The technique was first applied to fair division by [91]; below we generalize his results.
Lemma 4.
For any allocation $\mathbf{z^{*}}$:
(a) There exists an allocation $\mathbf{z}$ with at most $n$ sharings in which each agent receives exactly the same utility: $\forall i\in[n]:u_{i}(\mathbf{z}_{i})=u_{i}(\mathbf{z^{*}}_{i})$. The $n$ is tight even for shared objects.
(b) There exists an allocation $\mathbf{z}$ with at most $n-1$ sharings in which each agent receives at least as much utility: $\forall i\in[n]:u_{i}(\mathbf{z}_{i})\geq u_{i}(\mathbf{z^{*}}_{i})$. The $n-1$ is tight even for shared objects.
Proof.
For each $i\in[n]$, define the constant $U_{i}:=u_{i}(z^{*}_{i})$.
(a) The allocation $\mathbf{z}$ can be found using the following LP, in which the variables are the $z_{io}$ for all $i\in[n],o\in[m]$, and the constraints are:
$\sum_{i=1}^{n}z_{io}=1$ for $\text{for }o\in[m]$ (2)
$\sum_{o=1}^{m}v_{io}z_{io}=U_{i}$ for $\text{for }i\in[n]$
$z_{io}\geq 0$ for $\text{for }i\in[n],o\in[m]$
This LP has at least one solution ($\mathbf{z^{*}}$) and has $m+n$ equational constraints, so it has a solution $\mathbf{z}$ in which at most $m+n$ variables are non-zero: $\sum_{o\in[m]}\left(\big{|}\{i\in[n]:\,z_{i,o}>0\}\big{|}\right)\leq m+n$. By definition of sharing, the number of sharings is smaller or equal than $(m+n)-m=n$.
For tightness, consider an instance with $n$ agents and $n$ goods, where agent $i$ values good $i$ at $n^{2}-n+1$ and the other $n-1$ goods at $1$. Let $\mathbf{z}^{*}$ be the allocation giving each agent $1/n$ of each good. This allocation has $n$ sharings and $n$ shared objects, and gives each agent a utility of exactly $n$. But no allocation with less than $n$ shared objects can attain exactly the same utilities: if object $i$ is given entirely to agent $i$, then agent $i$’s utility is more than $n$; if it is given entirely to another agent, then agent $i$’s utility is at most $n-1$.
(b) The above LP can be modified by removing the constraint for agent $n$, and maximizing agent $n$’s utility instead:
maximize $\sum_{o=1}^{m}v_{no}z_{no}$ (3)
subject to $\sum_{i=1}^{n}z_{io}=1$ for $\text{for }o\in[m]$
$\sum_{o=1}^{m}v_{io}z_{io}=U_{i}$ for $\text{for }i\in[n-1]$
$z_{io}\geq 0$ for $\text{for }i\in[n],o\in[m]$
There is a feasible solution ($\mathbf{z^{*}}$) in which the objective value is $U_{n}$, so in every optimal solution the utility of agent $n$ is at least $U_{n}$ and the utility of each agent $i<n$ is exactly $U_{i}$. The LP has $m+n-1$ equational constraints, so it has a solution $\mathbf{z}$ with at most $m+n-1$ non-zeros. In this solution the number of sharings is smaller or equal than $(m+n-1)-m=n-1$.
For tightness, consider an instance with $n$ agents and $n-1$ goods whose value for all agents is $n$. Let $\mathbf{z^{*}}$ be the allocation that gives each agent $1/n$ of each good. This allocation has $n-1$ sharings and gives each agent a utility of $n-1$. But in any allocation with less than $n-1$ shared objects, at least one object is given entirely to some agent, and thus the remaining $n-1$ agents must share a total value of at most $(n-2)n<(n-1)(n-1)$; thus at least one remaining agent gets less than $(n-1)$. ∎
Note that Lemma 4 holds whether the valuations are positive, negative or mixed.
As a corollary, we two upper bounds (some of them already proved by [82]).
Theorem 0.B.1.
Any instance with $n$ agents admits a proportional allocation with at most $n-1$ sharings. The $n-1$ is tight even for shared objects.
Proof.
Apply Lemma 4(b) with $\mathbf{z^{*}}$ the allocation giving each agent $1/n$ of each object. For tightness, consider $n-1$ objects and $n$ agents who value all objects at $1$. A proportional allocation requires each agent to get exactly $(n-1)/n$ of an object, so all objects must be shared. ∎
Theorem 0.B.2.
Any instance with $n$ agents admits an envy-free allocation with at most $n-1$ sharings. The $n-1$ is tight even for shared objects.
Proof.
[26] prove the existence of a competitive equilibrium (CE) in the so-called Fisher market with equal budgets associated with the division problem, when valuations are additive but may have different signs. Equilibrium allocations satisfy the property of Pareto-indifference: if $\mathbf{z}^{*}$ is an equilibrium allocation and $\mathbf{z}$ gives the same utilities to all agents, then $\mathbf{z}$ is an equilibrium as well. It is known that competitive equilibrium allocations are envy-free, as well as fractionally Pareto-optimal. This allows us to apply Lemma 4(b) and get an envy-free allocation with at most $n-1$ sharings. The proof of tightness is the same as in (a). ∎
The following theorem was proved in an unpublished manuscript which is no longer available [91]; we repeat the proof briefly for completeness.
Theorem 0.B.3.
Any instance with $n$ agents admits an equitable allocation with at most $n-1$ sharings. The $n-1$ is tight even for shared objects.
Proof.
The following LP finds an equitable allocation which, moreover, maximizes the equal value (such an allocation is often called max-equitable):
maximize $\sum_{o=1}^{m}v_{no}z_{no}$ (4)
subject to $\sum_{i=1}^{n}z_{io}=1$ for $\text{for }o\in[m]$
${\sum_{o=1}^{m}v_{io}z_{io}\over V_{i}}={\sum_{o=1}^{m}v_{no}z_{no}\over V_{n}}$ for $\text{for }i\in[n-1]$
$z_{io}\geq 0$ for $\text{for }i\in[n],o\in[m]$
It has $m+n-1$ equational constraints; hence an equitable allocation with $n-1$ sharings exists. Tightness can be shown using the same argument as in Lemma 4(b)
The following result was proved constructively by [54]. Note that it does not follow e.g. from theorems in [25, 14, 82], since those theorems Pareto-improve the original allocation, while the consensus condition is not preserved under Pareto-improvements.
Theorem 0.B.4.
For any positive integers $n,k$, any instance with $n$ agents admits a $k$-consensus partition with at most $n(k-1)$ sharings. The $n(k-1)$ is tight even for shared objects.
Proof.
The following LP finds a $k$-consensus partition:
$\sum_{i=1}^{n}z_{io}=1$ for $\text{for }o\in[m]$ (5)
$\sum_{o\in[m]}v_{i,o}z_{j,o}=V_{i}/k$ for $\text{for }i\in[n],j\in[k-1]$
$z_{io}\geq 0$ for $\text{for }i\in[n],o\in[m]$
Note that for $j=k$ the equality holds automatically. Since there are $m+n(k-1)$ constraints, there exists a basic feasible solution with at most $n(k-1)$ sharings.
For tightness, consider an instance with $n$ agents and $n(k-1)$ goods. For each agent, $(k-1)$ of the goods are “big” and the other $(n-1)(k-1)$ goods are “small”. All $n$ sets of big goods are pairwise-disjoint, so that each good is “big” for exactly one agent and “small” for all other agents.
Each agent values each of his/her big goods at $k-0.5$ and each of his/her small goods at $0.5/(n-1)$. For each agent, the sum of all values is $k(k-1)$, so in a $k$-consensus partition, the value of each bundle should be exactly $k-1$. However, the value of each big good to its agent is larger than $k-1$, so all goods must be shared. ∎
Summarizing the above theorems, we have the following worst-case upper bounds on the number of sharings:
$S_{PROP}(n)=S_{EF}(n)=S_{EQ}(n)=n-1$ $S_{CONS}(n)=n(n-1)$
0.B.2 Computing an allocation attaining the worst-case upper bound
The previous theorems raise the computational question of how to find an allocation with at most $S_{F}(n)$ sharings given a fairness criterion F. The simplex method finds such solutions for any LP, and it may be sufficient for all practical purposes, but its worst-case run-time is exponential. [64] showed that the ellipsoid method can be used to find an optimal basic feasible solution in weakly-polynomial time, but its practical performance is not very good. Interior-point methods perform better both theoretically and practically, but they may find an interior solution (with many nonzeros) rather than a basic feasible solution. [73] gives a strongly-polynomial time algorithm that, given a pair of optimal solutions for the primal and dual problems, finds an optimal basic feasible solution. [22] give a different algorithm for the same task, that performs well in practice but does not have good worst-case run-time guarantees.
All the above algorithms still require a solution to the original LP. Here the fairness criteria differ greatly in their computational complexity.
For proportionality, it is sufficient to find an optimal solutions to (3) and its dual where $U_{i}=V_{i}/n$ for all $i\in[n]$; this can be done in weakly-polynomial time. Moreover, a proportional allocation with $n-1$ sharings can be found in strongly-polynomial time by reduction to connected cake-cutting in the following way. Arrange the objects on an interval in an arbitrary order, where each object corresponds to a homogeneous sub-interval; find a connected proportional division of the interval; and allocate the objects accordingly. Since a connected partition makes $n-1$ cuts, at most $n-1$ objects are shared. When all valuations are positive or all valuations are negative, a connected proportional allocation of the interval can be found using $O(mn\log(n))$ operations by the Even-Paz protocol [49]. When the valuations are mixed, a connected proportional allocation can be found using $O(mn^{2})$ operations by the Last Diminisher protocol [89].
For consensus, [54] present a polynomial-time algorithm that finds a $k$-consensus partition with at most $n(k-1)$ sharings. The main step in their algorithm is Gaussian elimination, which can be done in strongly-polynomial time; hence, their algorithm is in fact strongly-polynomial.
For envy-freeness, applying the proof of Theorem 0.B.2 requires to find a competitive equilibrium (CE) in a Fisher market with equal budgets. With strictly-positive valuations, a CE can be found in $O((n+m)^{4}\log(n+m))$ operations [79]. Note that reduction to connected cake-cutting (as we did for proportionality) is not very helpful with envy-freeness. The reduction yields a piecewise-homogeneous cake; for such cakes, the fastest algorithm that we know of for finding a connected EF allocation requires $O(m^{n})$ operations [3]. We do not know whether polynomial-time algorithms exist.
For equitability, optimal solutions to (4) and its dual can be found in weakly-polynomial time; we do not know if a strongly-polynomial time algorithm exists.
Open problem 5.
Can an equitable allocation with $\leq n-1$ sharings be found in strongly-polynomial time?
0.B.3 Hardness of finding an LP solution with fewer nonzero variables
As the number of sharings in an allocation is related to the number of nonzero variables in a solution of an LP, it could be very helpful to have an algorithm for finding a solution minimizing the number of nonzero variables in a given LP. This problem is known as Min-RVLS — Minimum Relevant Variables in a Linear System. Unfortunately, Min-RVLS is NP-hard, and it is also hard to find multiplicative approximations to the minimum [4]. Moreover, even finding a very modest additive approximation is NP-hard:
Theorem 0.B.5.
Unless P=NP, there is no polynomial-time algorithm that accepts as input a linear program with $k$ constraints and decides whether it has a feasible solution with at most $k-1$ nonzero variables.
Proof.
The proof is by reduction from Partition. Suppose we are given $k$ numbers $a_{1},\ldots,a_{k}$ whose total sum is $2S$, and have to decide whether they can be partitioned into two subsets such that the sum in each subset is $s$. This problem can be solves using an LP with $2k$ variables. For each $i\in[2]$ and $j\in[k]$, the variable $x_{ij}$ determines what fraction of the number $a_{j}$ is in subset $i$:
[…]
$\sum_{j=1}^{m}x_{1j}a_{j}=\sum_{j=1}^{m}x_{2j}a_{j}$
[…]
There are $k+1$ equational constraints, so there always exists a solution with at most $k+1$ nonzeros. In this solution, at most one number is “cut” between the sets. Indeed, it is easy to solve the partition problem if we are allowed to cut one number: just order the numbers on a line and cut the line into two parts with equal sum.
Now, if we could solve the “few nonzeros” problem, then we could decide whether the above LP has a solution with at most $k$ nonzeros, which would imply a solution to the partition problem (in which no number is cut). ∎
Open problem 6.
Is the problem of Theorem 0.B.5 strongly NP-hard?
Appendix 0.C Truthful Fair Division
A division algorithm is truthful if for every agent $i$, the utility of $i$ is maximized when $i$ reports the true valuations $(v_{i,o})_{o\in[m]}$. In general, truthfulness, fairness, and Pareto-optimality are incompatible, see [92]. However, truthfulness can be achieved by introducing some inefficiencies. This section surveys several such truthful mechanisms and checks whether any of them can be adapted to minimize the sharing.
0.C.1 Consensus allocation mechanism
[76] presents a truthful randomized mechanism that uses a consensus allocation. Given a consensus allocation $\mathbf{z}$, a permutation $\pi$ over $[n]$ is selected uniformly at random, and the bundle $\mathbf{z}_{\pi(i)}$ is allocated to agent $i$. Thus, the expected utility of any agent $i$, whether truthful or not, is $V_{i}/n$, so the agent cannot gain by false reporting. Moreover, a truthful agent gets a utility of exactly $V_{i}/n$ with certainty, while a non-truthful agent might get more or less than $V_{i}/n$. So for a risk-averse agent, truthfulness is a strictly dominant strategy.
Combining the algorithm of [76] with Theorem 0.B.4 gives:
Corollary 1.
There exists a randomized truthful algorithm that (with certainty) returns an envy-free allocation with at most $n(n-1)$ sharings.
The $n(n-1)$ is tight for a consensus allocation, but not necessarily tight for truthfulness. Below we survey some other approaches for truthful fair allocation.
0.C.2 Partial-allocation and strong-demand mechanisms
[44] suggest a different approach to truthful fair division, called Partial Allocation Mechanism (PAM). Their benchmark for fairness is the max-product allocation — the allocation maximizing the product of utilities (also known as the Nash-optimal or the Proportionally-fair allocation). Informally, PAM works as follows.
  1. 1.
    Find a max-product allocation $\mathbf{z^{*}}$.
  2. 2.
    For each agent $i$, compute the ratio $f_{i}$ between the product of the other agents’ utilities when $i$ is present, and the maximum product of their utilities when $i$ is not present.
  3. 3.
    Give to each agent $i$ the bundle $f_{i}\cdot\mathbf{z^{*}_{i}}$ (that is, a fraction $f_{i}$ of each good $i$ receives in the original max-product allocation).
They prove that, under reasonable assumptions, $f_{i}\geq 1/e\approx 0.368$, so each agent is guaranteed at least $36.8\%$ of his/her max-product utility. Moreover, they prove that, with additive linear utilities, the allocation is envy-free. It remains to analyze how many sharings this allocation requires.
It is well-known that, when all valuations are linear and additive as in our case, the max-product allocation is equivalent to the Competitive Equilibrium from Equal Incomes (CEEI) allocation which is a well-known rule for fair allocation of resources among agents with different preferences. As explained in Section 0.B, there always exists a CEEI allocation with at most $n-1$ sharings. However, the process of giving each agent $i$ only a fraction $f_{i}$ of each object is, arguably, equivalent to forcing the agent to share each and every object (e.g. with the mechanism designer, the government or the public).
A possible way to overcome this issue is to interpret $f_{i}$ not as a fraction but as a probability, that is: with probability $f_{i}$, agent $i$ receives his/her share in the max-product allocation, otherwise agent $i$ receives nothing. Since the utilities of all agents (assuming they are risk-neutral) are the same in both cases, the randomized mechanism is still truthful and has at most $n-1$ sharings. However, in contrast to the randomized mechanism based on consensus allocation, it does not guarantee proportionality ex-post, nor even ex-ante (since $1/e$ of the max-product allocation may be less than $1/n$ of the total value). [44] presents a different mechanism called Strong Demand Mechanism (SDM), which is particularly efficient when there are many agents and few goods (as in the case of coupon-based privatization). SDM gives to each agent a fraction of a single good, so all objects are shared. Hence, SDM is apparently incompatible with sharing minimization.
Remark 1.
For the case of two agents, [45] present a mechanism called MAX, which runs both the consensus-allocation mechanism and the partial-allocation mechanism and returns the outcome with the highest social welfare. They prove that the resulting mechanism is truthful, envy-free, proportional, and guarantees at least $2/3$ of the maximum social welfare. Moreover, as explained above, for two agents the consensus-allocation mechanism can be implemented with at most 2 sharings and the partial-allocation mechanism can be implemented with at most 1 sharing.
0.C.3 Mechanism for binary valuations
In the problem of fair cake-cutting, most positive results for truthful allocation are designed for the special case in which the agents have piecewise-uniform valuations. the cake is assumed to be an interval; each agent $i$ desires a finite set of sub-intervals of the cake, and does not care about the rest of the cake. For this setting, [43] presents a truthful mechanism, called the CLPP mechanism, that guarantees envy-freeness and Pareto-optimality. Informally, the CLPP mechanism proceeds as follows.
  1. 1.
    For each subset $S$ of agents, calculate their average size — the total size of intervals desired by at least one member of $S$.
  2. 2.
    Pick an $S$ with a smallest average size (breaking ties arbitrarily);
  3. 3.
    Allocate to the members of $S$, all their desired sub-intervals, such that each member of $S$ gets a value of exactly the average size (this means that all the desired sub-intervals are allocated to agents who desire them).
  4. 4.
    Divide the remaining cake recursively among the remaining agents.
  5. 5.
    Discard all parts of the cake that are undesired by any agent.
In our setting, piecewise-uniform valuations means that each agent assigns to each object a value of either $1$ or 00. Under this assumption, an agent cares only about the total amount he/she gets from desired objects. Therefore, in step 3, the allocation can be implemented using at most $|S|-1$ sharings. When $S$ is the set of all agents, the allocation requires $n-1$ sharings, which is the worst-case upper bound on any fair division.
Corollary 2.
For binary valuations, there exists a deterministic truthful algorithm that (with certainty) returns an envy-free and Pareto-optimal allocation with at most $n-1$ sharings.
Open problem 7.
(a) In some situations, it may be possible to attain the same allocation as the CLPP mechanism with less than $n-1$ sharings. Is it possible to minimize the amount of sharing, while keeping the mechanism truthful, envy-free and Pareto-optimal?
(b) What is the smallest number of sharings required for a truthful mechanism for finding proportional / envy-free / equitable allocation?
0.C.4 Truthfulness with mixed divisible and indivisible goods
[66] study the setting in which only some of the items, which are pre-determined as divisible, may be shared. Their goal is to design truthful mechanisms that also attain an approximation of envy-freeness called EFM (envy-freeness for mixed goods). They show that, in general, no truthful EFM algorithm exists, even if there is only one indivisible good and one divisible good and only two agents. But, when agents have binary valuations on indivisible goods and identical valuations on a single divisible good, an EFM and truthful mechanism exists. When agents have binary valuations over both divisible and indivisible goods, an EFM and truthful mechanism exists when there are only two agents, or when there is a single divisible good. All these results are not applicable to our setting, in which all goods may be divided if needed.
Appendix 0.D Further Related Work
Most traditional works on fair division either assume that all objects are divisible, or that all objects are indivisible. But in recent years, there is a growing interest in “mixed” settings; see [69] for a survey. The works most closely related to ours are:
  • Bei, Li, Liu, Liu and Lu [16] and Bei, Liu, Lu and Wang [18] study an allocation problem where some goods are divisible and some are indivisible. In their setting the distinction between divisible and indivisible objects is given in advance — the algorithm can only divide objects that are pre-determined as divisible. This means that an exact-fair allocation may not be possible, so they study approximate-fairness notions.
  • Bhaskar, Sricharan and Vaish [19] study a mixture of indivisible and divisible chores (objects with negative utilities), and present an algorithm for approximate envy-free allocation.
  • Nishimura and Sumita [77] and Kawase, Nishimura and Sumita [61] study welfare-maximization problems in settings of mixed goods.
  • Bei, Liu and Lu [17] generalize the above model by allowing subjective divisibility, wherein each object can be considered divisible by some agents and indivisible by others. Again, exact-fair allocation may not be possible, but they show algorithms for approximate envy-free and maximin-share allocations, focusing mainly on two and three agents.
  • Li, Li, Liu and Wu [65] study a setting in which each agent may have a different “indivisibility ratio” (= proportion of items that are indivisible), and the approximate-fairness guarantee for each agent depends on his indivisibility ratio. This is another way to allow indivisible objects but constraining them. While they bound the number of divisible objects per agent, we study a more global bound in the number of objects.
  • Li, Liu, Lu, Tao and Tao [67] study the price of fairness in both indivisible and mixed item allocation.
A crucial difference between these works and our paper is that we insist on exact fairness. This is important in settings with highly valuable objects, in which approximate fairness may be unacceptable.
0.D.1 Fairness with indivisible objects.
The Adjusted Winner procedure was applied (at least theoretically) to division problems in divorce cases and international disputes [33, 71] and was studied empirically [83, 47].
With indivisible objects, envy-free and even proportional allocations may not exist. The most commonly studied relaxations of these two concepts are Envy-freeness except one good (EF1) and Maximin share guarantee (MMS).
EF1 was introduced by [35]; a closely related concept was presented earlier by [68]. EF1 is widely studied, for example by [2, 78] and others. A strengthening of EF1 to a global fairness notion, based on information withholding, was recently studied by [58]. Existence of EF1 indivisible Pareto Optimal allocations was proved by [41] and [15] strengthened the result to fractional PO.
[35] defined MMS, demonstrated existence under large-market assumption and applied the concept in practice for course allocation in [36]. For “small markets”, [81] showed that MMS allocations may not exist for some knife-edge instances and hence all the results about MMS consider a certain approximation to MMS itself, e.g.,  [5, 13, 53, 52, 9, 10, 86].
Classic microeconomics mostly works with divisible resources and handles indivisible objects by making them “divisible” via a lottery. This approach results in weaker fairness: an allocation is fair ex-ante, i.e., in expectation before the lottery is implemented. Ex-ante fairness was analyzed in many different contexts. Just to name a few: [59], [1], [23] considered the problem of fair assignment and [37] considered its multi-unit constrained modifications; [63] evaluated fairness of tie-breaking in matching markets; [24] studied randomized rules for online fair division.
[30] observed that exact envy-freeness with indivisible goods can be achieved by leaving some of them unallocated while keeping some efficiency guarantees: their AL procedure constructs an allocation that is not Pareto-dominated by another envy-free allocation, see also [6]. Recently, [39] showed that, by leaving some items unallocated, it is possible to construct an allocation that is “Envy-free except any good” (also known as EFX, an approximate fairness notion which strengthens EF1 [41]).
[56] suggested a novel approach to achieve exact fairness with indivisible goods: introduce a monetary subsidy by a third party, while minimizing the transfers. This approach is further developed in the follow-up papers [34, 40]. Minimization makes it methodologically similar to our approach and distinguishes it from other results on fair allocation of indivisible goods with monetary transfers, e.g. the rent-division problem of [50].
There is a significant gap between worst-case and average-case numbers of sharings; this further stresses the importance of sharing minimization. [48] considered random instances and demonstrated that the minimal number of sharings for an envy-free fPO allocation is zero with high probability, if the number of goods is large and values are independent and identically distributed. [70] extended the result to allocations among agent groups.
0.D.2 Checking existence of fair allocations of indivisible objects.
Fair allocation of indivisible goods might not exist in all cases, but may exist in some. A natural question is how to decide whether it exists in a given instance. It was studied by [68, 62, 28] for various fairness and efficiency notions, showing that it is computationally hard in general (with some exceptions). The undercut procedure of [29] finds an envy-free allocation of indivisible goods among two agents with monotone (not necessarily additive) valuations, if-and-only-if it exists (see also [7]).
0.D.3 Cake-cutting with few cuts.
The goal of minimizing the number of “cuts” has also been studied in the context of fair cake-cutting — dividing a heterogeneous continuous resource, see [90, 87, 11, 12, 3, 84, 85, 46]. Since the resource is continuous, the techniques and results are quite different.
Appendix 0.E Generic NP-hardness Proofs for Section 5
Proposition (Proposition 2).
If a problem has a generically-polynomial algorithm, then it is not generically-NP-hard unless P=NP.
Proof.
Let $P_{2}$ be a decision problem that satisfies the theorem assumptions:
(a) There is a generically-polynomial algorithm $A_{2}$ for $P_{2}$; denote by $f_{p}$ the polynomial function in Definition 2.
(b) There is a multi-reduction $h_{t_{1}}$ from some NP-hard problem $P_{1}$ to $P_{2}$; denote by $f_{e}$ the super-polynomial function in Definition 5.
Since $f_{p}$ is polynomial and $f_{e}$ is super-polynomial, there is some $t_{0}$ such that $f_{e}(t)>f_{p}(t)$ for all $t>t_{0}$. We show a polynomial-time algorithm for solving $P_{1}$ on all inputs of size $t_{1}$ such that $t_{2}(t_{1})>t_{0}$.
Given an input $x_{1}\in\mathcal{I}({t_{1},c_{1}})$ to problem $P_{1}$, we use the assumed multi-reduction $h_{t_{1}}$ to compute in polynomial time an input $x_{2}\in\mathcal{I}({t_{2},c_{2}})$ to problem $P_{2}$.
By Definition 2(a,c), it is possible to compute in time poly$(t_{2})$ another input $x_{2}^{\prime}\in\mathcal{B}(x_{2},f_{p}(t_{2}))$, such that $A_{2}$ runs on $x_{2}^{\prime}$ in time poly$(t_{2})$.
But by definition of multi-reduction (Definition 4), $P_{2}(x_{2}^{\prime\prime})=P_{1}(x_{1})$ for all $x_{2}^{\prime\prime}\in\mathcal{B}(x_{2},f_{e}(t_{2}))$. In particular, since $f_{e}(t_{2})>f_{p}(t_{2})$, this holds for $x_{2}^{\prime}$ too, so $P_{2}(x_{2}^{\prime})=P_{1}(x_{1})$.
Therefore, by running $A_{2}$ on $x_{2}^{\prime}$, we get the correct answer to $P_{1}(x_{1})$. ∎
Theorem (Theorem 5.1).
The decision problems Prop-Sharings$(n,s)$ and EF-Sharings$(n,s)$ are generically NP-hard in the following cases:
(a) For any fixed $n\geq 2$ and $s=0$;
(b) For any fixed $n\geq 3$ and $s\in\{0,\dots,n-3\}$.
Proof.
We show a multi-reduction from $P_{1}=$ $k$-way Partition to $P_{2}=$ F-Sharings$(n,s)$.
The input to $P_{1}$ is a list $D:=[d_{1},\ldots,d_{t_{1}}]$ of integers in $[0,2^{c_{1}t_{1}}]$, where $c_{1}\geq 1$ is some integer constant.666 In fact, we can assume $c_{1}=1$. The NP-hardness proof of Partition described in [51] constructs instances with some $p$ integers, in which each input integer can be represented by at most $3p\log(p)$ bits. We can add to each constructed instance $3p\log(p)-p$ zeros; this does not affect the existence of a partition. Then, taking $t_{1}:=3p\log(p)$, we have that each constructed instance has $t_{1}$ integers in $[0,2^{t_{1}}]$. So Partition is NP-hard even when restricted to $\mathcal{I}({t,c})$ with $c=1$. The task is to decide if the integers can be partitioned into $k$ subsets with sum $S:=\frac{1}{k}\sum_{o=1}^{t_{1}}d_{o}$. We denote $L:=2^{c_{1}t_{1}}$.
For (a), we reduce from $k$-way Partition for $k=n$. We construct an instance of F-Sharings$(n,0)$ with $m=t_{1}+n$ objects. Note that the number $t_{2}$ of inputs to $P_{2}$ is the number of values in the valuation matrix, $t_{2}=mn=nt_{1}+n^{2}$, which is polynomial in $t_{1}$ as $n$ is fixed.
All agents’ valuations will be integers in the range $[0,2^{c_{2}t_{2}}]$, where $c_{2}:=2c_{1}$. For simplicity, we denote by “///” the integer division operation.
  • The objects $o\in\{1,\ldots,t_{1}\}$ are usual objects. Each agent $i\in[n]$ values each usual object $o$ at $L\cdot d_{o}+L/(16m)$.
  • The objects $o\in\{t_{1}+1,\ldots,t_{1}+n\}$ are compensation objects. Each agent $i\in[n]$ values the compensation object $t_{1}+i$ at $83L/$ 8, and every other compensation object at $L/(16mn)$.
Note that all valuations are integers smaller than $L^{2}+L$, which is indeed smaller than $2^{c_{2}t_{2}}$.
For the multi-reduction we will use the function
$f_{e}(t_{2})$ $:=L/(64mn)=2^{t_{1}c_{1}-6}/(mn)$
$=2^{c_{1}(t_{2}-n^{2})/n}/t_{2}.$
Note that $f_{e}$ is indeed a super-polynomial function of $t_{2}$.
Suppose that there exists a partition of $D$ into $n$ subsets $(D_{1},\ldots,D_{n})$ with sums equal to $S$. We construct an allocation of the objects by giving each agent $i$ the set of usual objects corresponding to the integers in $D_{i}$, as well as the compensation object indexed $t_{1}+i$. We show that the resulting allocation is EF (and hence PROP) not only for the constructed valuation matrix, but also for any valuation matrix with infinity-distance of at most $L/(64mn)$.
Indeed, in every such input, each agent values usual object $o$ at least $L\cdot d_{o}+L/(16m)-L/(64mn)>L\cdot d_{o}$, and values his compensation object at least $83L/8-L/(64mn)>2L/$ 8, so he values his own bundle at least $LS+2L/8$.
On the other hand, each agent values usual object $o$ at most $L\cdot d_{o}+L/(16m)+L/(64mn)<L\cdot d_{o}+L/(8m)$, and values any other compensation object at most $L/(16mn)+L/(64mn)<L/(8mn)$. Therefore, each agent values the bundle of any other agent at most $LS+m\cdot(L/8m)=LS+L/8$, so there is no envy.
Conversely, suppose there is a PROP allocation (or an EF allocation, which is always PROP). As each agent values each usual object $o$ at least $Ld_{o}$, the sum of all values for the agent is larger than $L\cdot nS$, so PROP requires each agent to receive a value larger than $LS$.
The value of each potential bundle comes from integer multiples of $L$ (— the $L\cdot d_{o}$ value of usual objects), and fractions of $L$ (— the $L/(16m)$ value of usual objects, and the values of compensation objects). The sum of all fractions is strictly less than $L$. Therefore, to get a value of at least $LS$, each agent must receive usual objects with $\sum_{o}d_{o}\geq S$. The partition of usual objects corresponds to a partition of $D$ into $n$ subsets of sum at least $S$ each. But the total sum of all integers is $nS$, so the sums of all $n$ subsets must be exactly $S$.
For (b), we reduce from $k$-way Partition for $k=n-s-1$. Note that $k\geq 2$ since $s\leq n-3$. We construct an instance of F-Sharings$(n,s)$ with $n$ agents, of whom $k$ are usual agents constructed similarly to the $k$ agents in part (a), and the remaining $s+1$ are special agents, indexed $k+1,\ldots,k+s+1$. As in part (a), there are $t_{1}$ usual objects and $n$ compensation objects. In addition, there is a single special object. All in all we have $n=k+s+1$ and $m=t_{1}+n+1$. The additional valuations are:
  • Each usual agent values the special object at $(s+1)LS+L/16$
  • Each special agent $i$ values the special object at $n\cdot LS+L/16$, and values each of the other objects (usual and compensation) at $L/(16mn)$.
We use the same super-polynomial function $f_{e}(t_{2})=L/(64mn)$ as in part (a).
Suppose there exists a partition of $D$ into $k$ subsets $(D_{1},\ldots,D_{k})$ with sums equal to $S$. Similarly to part (a), we give each usual agent $i$ the set of usual objects corresponding to the integers in $D_{i}$, as well as the compensation object indexed $t_{1}+i$. We divide the special object equally among the $s+1$ special agents, so that there are $s$ sharings. We show that the resulting allocation is EF (and hence PROP) not only for the constructed valuation matrix, but also for any valuation matrix with infinity-distance of at most $L/(64mn)$.
As in part (a), each usual agent $i\in\{1,\ldots,k\}$ values his own bundle at least $LS+2L/8$, and values the bundle of every other usual agent at most $LS+L/8$. Moreover, each usual agent values the special object at most $LS+L/16+L/(64mn)<LS+L/8$, so values the bundle of every special agent at most $LS+(L/8)/(s+1)$, so the usual agents do not envy.
Each special agent $i\in\{k+1,\ldots,k+s+1\}$ values the special object at least $nLS+L/16-L/(64mn)>nLS$, so values his own bundle at least $(nLS)/(s+1)>LS$. Moreover, each special agent values every other object at most $L/(16mn)+L/(64mn)<L/(8mn)$, so values every other bundle at most $L/8$. Therefore, the special agents do not envy too, so the allocation is EF and hence also PROP.
Conversely, suppose there is a PROP allocation (or an EF allocation, which is PROP) with at most $s$ sharings. Then:
- Each special agent must receive a value of at least $(nLS)/n=LS$. Since their total value for all objects except the special object is at most $L/8$, each special agent must receive a part of the special object. This means that all $s$ sharings are in the special object, so the usual and compensation objects must be allocated without sharing.
- For each usual agent, the sum of all values is larger than $kLS+(s+1)LS=nLS$, so they must receive a value larger than $LS$ each. As in part (a), the sum of all fractions in the values of usual and compensation objects is smaller than $L$. Furthermore, no more sharings are allowed in the special object. So each usual agent must receive usual objects with $\sum_{o}Ld_{o}\geq LS$, and this induces a $k$-partition of $D$ into $k$ subsets of sum exactly $S$. ∎
We could not adapt the proof of Theorem 5.1 to shared objects. We could potentially have $s$ special objects that have to be shared, but then some parts of the special objects could benefit the usual agents, which would undermine the claim in the last paragraph of that proof. We can still prove similar results for shared objects, but with a different multi-reduction.
Theorem (Theorem 5.2).
The problems Prop-SharedObj$(n,s)$ and EF-SharedObj$(n,s)$ are generically NP-hard in the following cases:
(a) For any fixed $n\geq 2$ and $s=0$;
(b) For any fixed $n\geq 3$ and $s\in\{0,\dots,n-3\}$.
Proof.
We prove both parts using a single multi-reduction from the case of identical valuations, proved to be NP-hard in [21].
We are given a list $D$ of items777we use “items” for the original instance $D$, and “objects” for the constructed fair allocation instance. whose values are integers $[d_{1},\ldots,d_{t_{1}}]$ in $[0,2^{c_{1}t_{1}}]$, where $c_{1}\geq 1$ is some constant, with $\sum_{o=1}^{t_{1}}d_{o}=nS$. We have to decide if they can be partitioned into $n$ bins with sum $S$ each, with at most $s$ items split between two or more bins. Let $L:=2^{c_{1}t_{1}}$.
We construct a fair allocation instance with $n$ agents and $m=t_{1}+n$ objects. Hence, $t_{2}=mn=nt_{1}+n^{2}$, which is linear in $t_{1}$ as $n$ is fixed. All agents’ valuations will be integers in $[0,2^{c_{2}t_{2}}]$, where $c_{2}:=2c_{1}$. Recall that “///” denotes integer division.
  • The objects $o\in\{1,\ldots,t_{1}\}$ are usual objects. Each agent $i\in[n]$ values each usual object $o$ at $L\cdot d_{o}+L/(16mn)$.
  • The objects $o\in\{t_{1}+1,\ldots,t_{1}+n\}$ are compensation objects. Each agent $i\in[n]$ values the compensation object $t_{1}+i$ at $3L/(8n)$ and every other compensation object at $L/(16mn)$.
Note that all valuations are integers smaller than $L^{2}+L$, which is indeed smaller than $2^{c_{2}t_{2}}$. For the multi-reduction we use the same super-polynomial function as in Theorem 5.1: $f_{e}(t_{2}):=L/(64mn)$.
Suppose that there exists a partition of $D$ into $n$ subsets $(D_{1},\ldots,D_{n})$ with sums equal to $S$, with at most $s$ shared items. We construct an allocation of the objects in our instance by giving each agent $i$ the set of usual objects corresponding to the items in $D_{i}$ (including fractions), as well as the compensation object $t_{1}+i$. Note that the number of shared items remains unchanged (at most $s$).
We now prove that the allocation is EF (and hence PROP) not only for the constructed instance, but also for any instance with infinity distance at most $L/(64mn)$. In all such instances, each agent values every usual object $o$ at least $Ld_{o}+L/(16mn)-L/(64mn)>Ld_{o}$ and his compensation object at least $3L/(8n)-L/(64mn)>2L/(8n)$, so values his own bundle at least $LS+2L/(8n)$; and values the bundle of every other agent at most $S+L/(8n)$. Hence the allocation is EF and also PROP.
Conversely, suppose there is a PROP allocation (or an EF allocation, which is always PROP), with at most $s$ shared objects. For each agent $i$, the sum of all object values is larger than $L\cdot nS$, so PROP requires each agent to receive a value larger than $LS$.
Denote by $D_{i}$ the set of items from $D$ (including fractions) corresponding to the usual objects given to agent $i$.
The value of each potential bundle comes from integer multiples of $L$ (— the $L\cdot d_{o}$ value of usual objects), and fractions of $L$ (— the $L/(16mn)$ value of usual objects, and the values of compensation objects). The sum of all fractions is strictly less than $L/n$. Therefore, to get a value of at least $LS$, each agent must receive usual objects with $\sum_{o\in D_{i}}Ld_{o}>LS-L/n$, which implies $\sum_{o\in D_{i}}d_{o}>S-1/n$ for each agent $i\in[n]$. Since $\sum_{o\in D}=nS$, this implies $\sum_{o\in D_{j}}d_{o}<S+1$ for each agent $j\in[n]$. Therefore, each bundle $D_{i}$ with $\sum_{o\in D_{i}}d_{o}\neq S$ must have fractions of items (as all item values in $D$ are integers). Therefore, we can move fractions of items from bundles $D_{i}$ with sum larger than $S$ to bundles $D_{j}$ with sum smaller than $S$, without increasing the number of shared items. By iteratively moving these fractions, we can construct in polynomial time a partition of $D$ with at most $s$ shared items, in which all bin sums equal $S$. ∎
We could not adapt the proof of Theorem 5.2 to sharings, since moving fractions of items between bins might increase the number of sharings. This is why we used a different proof for sharings in Theorem 5.1.
Theorem (Theorem 5.3).
For any fixed integer $n\geq 2$, the decision problems PropDpo-Sharings$(n,0)$ and EFDpo-Sharings$(n,0)$ are generically NP-hard.
Proof.
We adapt the proof of Theorem 5.1(a). It is sufficient to prove that, if there exists an equal-sum partition of $D$, then there exists a dPO+EF allocation (with no sharings) of the objects.
Let $\mathbf{z}$ be the EF allocation constructed in the proof of Theorem 5.1(a). If $\mathbf{z}$ is also dPO then we are done. Otherwise, let $\mathbf{y}$ be a dPO allocation (with no sharing) that Pareto-dominates $\mathbf{z}$. We claim that $\mathbf{y}$ is EF too.
Since $\mathbf{y}$ dominates $\mathbf{z}$, each agent must receive in $\mathbf{y}$ a value of at least $LS+2L/8>LS$. As in the proofs of the previous theorems, this means that each agent must receive some usual objects corresponding to integers from $D$ with a sum at least $S$. But since the sum of all integers in $D$ is $nS$, each agent must receive usual objects corresponding to integers of sum exactly $S$. This means that each agent $i$ values the bundle of every other agent at most $LS+L/8$. Therefore, the allocation $\mathbf{y}$ is EF, so it is the desired dPO+EF allocation. ∎
Appendix 0.F Strong generic NP-hardness
In many practical problems on integers, it can be assumed that the input integers are not too large; specifically, that the magnitude of each input integer is polynomial (rather than exponential) in the number of inputs. Equivalently, the binary encoding length of each input integer is logarithmic (rather than polynomial) in the number if inputs.
Formally, denote by $\mathcal{I}_{p}({t,c})$ the set of $t$-sized vectors of integers in $[0,2^{c\cdot\log_{2}t}]=[0,t^{c}]$ for some constant $c$. The following two classic concepts are used to describe run-time complexity of problems restricted to inputs in $\mathcal{I}_{p}({t,c})$:
  • An algorithm is said to run in pseudo-polynomial-time if it runs in time poly$(t)$ when restricted to inputs from $\mathcal{I}_{p}({t,c})$;
  • An algorithmic problem on integers is said to be strongly NP-hard if it is NP-hard even when restricted to inputs from $\mathcal{I}_{p}({t,c})$.
It is well-known that, for any fixed $n$, the problem $n$-way Partition has a pseudo-polynomial-time algorithm (based on dynamic programming). Similarly, the F-SharedObj$(n,s)$ and F-Sharings$(n,s)$ can be solved in pseudo-polynomial time for any fixed $n$ using dynamic programming. In contrast, when $n$ is part of the input, these problems are strongly NP-hard, which means that they do not have a pseudo-polynomial-time algorithm unless P=NP.
We would like to extend the notions of generically-easy and generically-hard to problems restricted to $\mathcal{I}_{p}({t,c})$. The definitions below are almost identical to the analogous definitions from Section 5, except that $t$ is replaced with $\log_{2}t$.
The following is analogous to Definition 2:
Definition 6.
Given an algorithm $A$ for a problem $P$, we say that $A$ runs generically in pseudo-polynomial time if there exists a polynomial function $f_{p}$ such that, for every size $t$ and input $x\in\mathcal{I}_{p}({t,c})$, there is a subset of “Good inputs” $G(x)\subseteq\mathcal{B}(x,f_{p}(\log_{2}t))$, such that the following holds:
(a) $A$ runs in time poly($t$) on all inputs in $G(x)$;
(b) The fraction of good inputs approaches 1, that is,
$\lim_{t\to\infty}\min_{x\in\mathcal{I}({t,c})}\frac{|G(x)|}{|\mathcal{B}(x,f_{p}(\log_{2}t))|}=1$
(c) Given $x$, it is possible to compute in time poly($t$) a vector in $G(x)$.
The following is analogous to Definition 4:
Definition 7 (strong multi-reduction).
Given two decision problems, $P_{1}$ defined on inputs in $\mathcal{I}({t_{1},c_{1}})$ for some constant $c_{1}\geq 1$ and $P_{2}$ defined on inputs in $\mathcal{I}({t_{2},c_{2}})$ for some constant $c_{2}\geq 1$, a strong multi-reduction from $P_{1}$ to $P_{2}$ consists of a polynomial-time computable function $t_{2}:\mathbb{N}\to\mathbb{N}$ and a family of functions $h_{t_{1}}:\mathcal{I}_{p}({t_{1},c_{1}})\to\mathcal{I}_{p}({t_{2}(t_{1}),c_{2}})$, which maps an input for $P_{1}$ to an input for $P_{2}$, and satisfies the following:
(a) $h_{t_{1}}$ runs in time poly$(t_{1})$;
(b) There exists an super-polynomial function $f_{e}$ such that, for all $t_{1}$ and all $x_{1}\in\mathcal{I}_{p}({t_{1},c_{1}})$, when $x_{2}:=h_{t_{1}}(x_{1})$,
[…]
Note the requirement that $h_{t_{1}}$ map each input to $\mathcal{I}_{p}({t_{1},c_{1}})$ rather than to $\mathcal{I}({t_{2},c_{2}})$; it implies that the reduction must keep the magnitude of the input integers polynomial in the number of inputs.
The following is analogous to Definition 5:
Definition 8.
A decision problem $P_{2}$ is called generically strongly NP hard if there exists a strong multi-reduction from some strongly NP-hard problem $P_{1}$ to $P_{2}$.
The following is analogous to the classic statement that a strongly NP-hard problem does not have a pseudo-polynomial-time algorithm unless P=NP. Its proof is almost identical to that of Proposition 2:
Proposition 3.
If a problem is generically-strongly-NP-hard, then it does not have a generically-pseudo-polynomial-time algorithm unless P=NP.
Proof.
Let $P_{2}$ be a decision problem that satisfies the theorem assumptions:
(a) There is a generically-pseudo-polynomial algorithm $A_{2}$ for $P_{2}$; denote by $f_{p}$ the polynomial function in Definition 6.
(b) There is a multi-reduction $h_{t_{1}}$ from some NP-hard problem $P_{1}$ to $P_{2}$; denote by $f_{e}$ the super-polynomial function in Definition 5.
Since $f_{p}$ is polynomial and $f_{e}$ is super-polynomial, there is some $t_{0}$ such that $f_{e}(\log_{2}t)>f_{p}(\log_{2}t)$ for all $t>t_{0}$. We show a polynomial-time algorithm for solving $P_{1}$ on all inputs of size $t_{1}$ such that $t_{2}(t_{1})>t_{0}$.
Given an input $x_{1}\in\mathcal{I}_{p}({t_{1},c_{1}})$ to problem $P_{1}$, we use the assumed multi-reduction $h_{t_{1}}$ to compute in polynomial time an input $x_{2}\in\mathcal{I}_{p}({t_{2},c_{2}})$ to problem $P_{2}$.
By Definition 6(a,c), it is possible to compute in time poly$(t_{2})$ another input $x_{2}^{\prime}\in\mathcal{B}(x_{2},f_{p}(\log_{2}t_{2}))$, such that $A_{2}$ runs on $x_{2}^{\prime}$ in time poly$(t_{2})$.
But by definition of multi-reduction (Definition 7), $P_{2}(x_{2}^{\prime\prime})=P_{1}(x_{1})$ for all $x_{2}^{\prime\prime}\in\mathcal{B}(x_{2},f_{e}(\log_{2}t_{2}))$. In particular, since $f_{e}(\log_{2}t_{2})>f_{p}(\log_{2}t_{2})$, this holds for $x_{2}^{\prime}$ too, so $P_{2}(x_{2}^{\prime})=P_{1}(x_{1})$.
Therefore, by running $A_{2}$ on $x_{2}^{\prime}$, we get the correct answer to $P_{1}(x_{1})$.
As $P_{1}$ is strongly-NP-hard, it is NP-hard even when restricted to inputs in $\mathcal{I}_{p}({t_{1},c_{1}})$. We have shown that $P_{1}$ can be solved in polynomial time for all inputs in $\mathcal{I}_{p}({t_{1},c_{1}})$ for sufficiently large $t_{1}$; hence P=NP. ∎
We are now ready to extend our hardness results to fair allocation problems in which $n$ is unbounded.
Theorem 0.F.1.
For every $s\geq 0$ Prop-Sharings$(s)$ and EF-Sharings$(s)$ are generically strongly NP-hard.
Proof.
We show a strong-multi-reduction from $P_{1}=$ 3-Partition, which is known to be strongly NP-hard. The input to $P_{1}$ is a finite multiset $D$ of $t_{1}=3p$ positive integers $d_{1},\ldots,d_{3p}$, each of which is in $[0,{t_{1}}^{6}]$.888 For the strong NP-hardness proof in [51], $c_{1}=6$ is sufficient. The sum of all integers is $pS$, for some $p\geq 2$ and integer $S$. The task is to decide if the integers can be partitioned into $p$ triplets such that the sum of each triplet is exactly $S$. We denote $L:={t_{1}}^{6}$. Note that we can assume w.l.o.g. that each integer in $D$ is in $(S/4,S/2)$, so every set with sum exactly $S$ must be a triplet.
We construct an instance of F-Sharings$(s)$ with $n=p+s+1$ agents and $m=3p+n+1=4p+s+2$ objects. Therefore, $t_{2}=mn=(p+s+1)(4p+s+2)$, which is in $O({t_{1}}^{2})$. The valuations are shown in the table below.
Agent $i$ / Object $o$: $o\in 1,\ldots,3p$ (“usual objects”) $o\in 3p+1,\ldots,3p+n$ (“compensation objects”) $o=3p+n+1$ (“special object”)
$i\in 1,\ldots,p$ (“usual agents”) $Ld_{o}+LS+L/(16m)$; $83L/$ 8 for $o=3p+i$; $L/(16mn)$ for $o\neq 3p+i$. $L\cdot(s+1)\cdot 4S+L/16$;
$i\in p+1,\ldots,n$ (“special agents”) $L/(16m)$; $L/(16m)$; $L\cdot n\cdot S+L/(16m)$;
All valuations are in $O(L^{2})=O({t_{1}}^{12})$, so we can let $c_{2}:=12$.
For the multi-reduction, we will use the function $f_{e}(\log t_{2})=L/(64mn)={t_{1}}^{6}/(64t_{2})\in\Omega({\sqrt{t_{2}}}^{6}/t_{2})=\Omega({t_{2}}^{2})$, which is indeed super-polynomial in $\log(t_{2})$.
Suppose $D$ has a partition into $p$ triplets, $D_{1},\ldots,D_{p}$, with sum $S$ each. We give each usual agent $i$ the three usual objects corresponding to the triplet $D_{i}$, and the compensation object $3p+i$. We divide the special object equally among the $s+1$ special agents, using $s$ sharings. We show that the resulting allocation is EF (and hence PROP) not only for the constructed valuation matrix, but also for any valuation matrix with infinity-distance of at most $L/(64mn)$. In all such instances:
- Each usual agent values usual object $o$ at least $Ld_{o}+LS+L/(16m)-L/(64mn)>Ld_{o}+LS$, and values his compensation object at least $83L/8-L/(64mn)>2L/$ 8, so he values his own bundle at least $LS+3LS+2L/8=4LS+2L/8$. On the other hand, each usual agent values usual object $o$ at most $Ld_{o}+LS+L/(16m)+L/(64mn)<Ld_{o}+LS+L/(8m)$, every other compensation object at most $L/(16mn)+L/(64mn)<L/(8mn)$, and the special object at most $L(s+1)\cdot 4S+L/8$, so he values the bundle of every other agent (usual or special) at most $84LS+L/$ 8. Therefore the usual agents do not envy.
- Each special agent values his own bundle, as well as all special agents’ bundles, at exactly the same value, which is at least $(nLS)/(s+1)>LS$. On the other hand, he values any non-special object at most $L/(8m)$, so values the bundles of usual agents at most $L/8$, so the special agents too do not envy.
Conversely, suppose there is a PROP allocation with at most $s$ sharings. Then:
- Each special agent must receive a value larger than $(nLS)/n=LS$. Since their total value for all objects except the special object is at most $L/8$, each special agent must receive a part of the special object. This means that all $s$ sharings are of the special object, so the usual and compensation objects must be allocated without sharing.
- Each usual agent values all objects together more than $4pLS+4(s+1)LS=4nLS$, so must receive a value larger than $4LS$. The value of each potential bundle without special objects comes from integer multiples of $L$ (— the $L\cdot d_{o}+LS$ value of usual objects), and fractions of $L$ (— the $L/(16m)$ value of usual objects, and the values of compensation objects). The sum of all these fractions is strictly less than $L$. Since no more sharings are allowed in the special, each usual agent must receive usual objects with no sharings, with total value at least $4LS$.
This induces a partition of $D$ into $p$ subsets with sum $\sum_{o}d_{o}\geq S$. As the sum of all $D$ is exactly $pS$, the sum of each part in the latter partition must be exactly $S$. As each integer in $D$ is in $(S/4,S/2)$, all parts must be triplets. ∎
Theorem 0.F.2.
For any fixed $s\geq 0$, Prop-SharedObj$(s)$ and EF-SharedObj$(s)$ are strongly NP-hard.
Proof.
The proof is very similar to that of Theorem 5.2, but adapted to a strong-multi-reduction from a strongly-NP-hard problem.
We show a multi-reduction from the case of identical valuations, proved to be strongly-NP-hard in [21].
An input of size $t_{1}$ consists of a list $D$ of $t_{1}-1$ items, whose values are integers $[d_{1},\ldots,d_{t_{1}-1}]$ in $[0,{t_{1}}^{c_{1}}]$, where $c_{1}$ is some constant, $c_{1}\geq 6$. The $t_{1}$-th input is $n$, which should be at most $t_{1}$. The sum of all values is $\sum_{o=1}^{t_{1}-1}d_{o}=nS$. We have to decide if they can be partitioned into $n$ bins with sum $S$ each, with at most $s$ items split between two or more bins. Let $L:={t_{1}}^{c_{1}}$.
We construct an instance of F-SharedObj$(s)$ with $n$ agents and $m=t_{1}-1+n$ objects. Hence, $t_{2}=mn+1=nt_{1}+n^{2}+1$, which is in $O({t_{1}}^{2})$. All agents’ valuations will be integers in $[0,{t_{2}}^{c_{2}}]$, where $c_{2}:=2c_{1}$. Recall that “///” denotes integer division.
  • The objects $o\in\{1,\ldots,t_{1}-1\}$ are usual objects. Each agent $i\in[n]$ values each usual object $o$ at $L\cdot d_{o}+L/(16mn)$.
  • The objects $o\in\{t_{1},\ldots,t_{1}-1+n\}$ are compensation objects. Each agent $i\in[n]$ values the compensation object $t_{1}-1+i$ at $3L/(8n)$ and every other compensation object at $L/(16mn)$.
Note that all valuations are integers smaller than $L^{2}+L$, which is indeed smaller than ${t_{2}}^{c_{2}}$. For the multi-reduction we use the function $f_{e}(\log t_{2}):=L/(64mn)={t_{1}}^{c_{1}}/(64(t_{2}-1))$. Note that $f_{e}(\log t_{2})\in\Omega(\sqrt{t_{2}}^{6}/t_{2})=\Omega({t_{2}}^{2})$, which is indeed super-polynomial in $\log t_{2}$.
Suppose that there exists a partition of $D$ into $n$ subsets $(D_{1},\ldots,D_{n})$ with sums equal to $S$, with at most $s$ shared items. We construct an allocation of the objects in our instance by giving each agent $i$ the set of usual objects corresponding to the items in $D_{i}$ (including fractions), as well as the compensation object $t_{1}-1+i$. The number of shared items remains unchanged, at most $s$.
We now prove that the allocation is EF (and hence PROP) not only for the constructed instance, but also for any instance with infinity distance at most $L/(64mn)$. In all such instances, each agent values every usual object $o$ at least $Ld_{o}+L/(16mn)-L/(64mn)>Ld_{o}$ and his compensation object at least $3L/(8n)-L/(64mn)>2L/(8n)$, so values his own bundle at least $LS+2L/(8n)$; and values the bundle of every other agent at most $S+L/(8n)$. Hence the allocation is EF and also PROP.
Conversely, suppose there is a PROP allocation (or an EF allocation, which is always PROP), with at most $s$ shared objects. For each agent $i$, the sum of all object values is larger than $L\cdot nS$, so PROP requires each agent to receive a value larger than $LS$.
Denote by $D_{i}$ the set of items from $D$ (including fractions) corresponding to the usual objects given to agent $i$.
The value of each potential bundle comes from integer multiples of $L$ (— the $L\cdot d_{o}$ value of usual objects), and fractions of $L$ (— the $L/(16mn)$ value of usual objects, and the values of compensation objects). The sum of all fractions is strictly less than $L/n$. Therefore, to get a value of at least $LS$, each agent must receive usual objects with $\sum_{o\in D_{i}}Ld_{o}>LS-L/n$, which implies $\sum_{o\in D_{i}}d_{o}>S-1/n$ for each agent $i\in[n]$. Since $\sum_{o\in D}=nS$, this implies $\sum_{o\in D_{j}}d_{o}<S+1$ for each agent $j\in[n]$. Therefore, each bundle $D_{i}$ with $\sum_{o\in D_{i}}d_{o}\neq S$ must have fractions of items (as all item values in $D$ are integers). Therefore, we can move fractions of items from bundles $D_{i}$ with sum larger than $S$ to bundles $D_{j}$ with sum smaller than $S$, without increasing the number of shared items. By iteratively moving these fractions, we can construct in polynomial time a partition of $D$ with at most $s$ shared items, in which all bin sums equal $S$. ∎